### v1

# Explainable Procedural Mistake Detection

Abstract
Automated task guidance has recently attracted attention from the AI research community. Procedural mistake detection (PMD) is a challenging sub-problem of classifying whether a human user (observed through egocentric video) has successfully executed the task at hand (specified by a procedural text). Despite significant efforts in building resources and models for PMD, machine performance remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we recast PMD to an explanatory self-dialog of questions and answers, which serve as evidence for a decision. As this reformulation enables an unprecedented transparency, we leverage a fine-tuned natural language inference (NLI) model to formulate two automated coherence metrics for generated explanations. Our results show that while open-source VLMs struggle with this task off-the-shelf, their accuracy, coherence, and dialog efficiency can be vastly improved by incorporating these coherence metrics into common inference and fine-tuning methods. Furthermore, our multi-faceted metrics can visualize common outcomes at a glance, highlighting areas for improvement.
Explainable Procedural Mistake Detection
Shane Storks        Itamar Bar-Yossef        Yayuan Li        Zheyuan Zhang        Jason J. Corso        Joyce Chai University of Michigan, Ann Arbor, Michigan, USA {sstorks, itamarby, yayuanli, zheyuan, jjcorso, chaijy}@umich.edu
Figure 1: Overview of self-reflective explanatory dialog between a foundational vision-and-language model (VLM) and itself to gather visual information and facilitate explainable procedural mistake detection (PMD).
1 Introduction
The problem of automated, interactive task guidance has recently attracted attention in the AI research community Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and interact with them through language to guide them through completing a task. One key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must apply physical commonsense reasoning to anticipate the success conditions for actions in this procedural text, then verify them by extracting relevant physical state information from the visual scene.
Prior work in PMD has explored two approaches: one thread of work has fine-tuned primarily vision-based classifiers without incorporating language Wang et al. (2023a); Peddi et al. (2024), while another has applied foundational language models (LMs) and vision-and-language models (VLMs) to this problem Du et al. (2023); Bao et al. (2023). However, both types of approaches have failed to achieve a viable level of accuracy in detecting mistakes. Qualitatively, Bao et al. (2023) find that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are generally unreliable procedural mistake detectors, transparency is crucial to ensure the human user can understand the system’s reasoning process and act on, disregard, or even correct its outputs accordingly. and the identification of areas for improvement.
To promote the development of PMD systems with greater transparency and coherence, we propose the reformulated problem of explainable PMD: given a procedural text and egocentric video frame, VLMs must not only decide whether a mistake has occurred, but additionally explain their decision with visual information from the scene, specifically through an iterative self-dialog of generated questions and answers which condition their decision. [^2] Of the many possible forms this explanation could take, we choose this one because of its resemblance to visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), tasks that have long been studied in the NLP and computer vision communities, and recent open-source VLMs are often trained and perform well on Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). We define two automated metrics for the coherence of generated explanations based on a fine-tuned natural language inference (NLI) model. To lay the groundwork for research into coherent explainable PMD, we apply three natural interventions to off-the-shelf VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness their in-context learning capability by prompting them with human-written examples to augment the space of candidate questions, and (3) we fine-tune VLMs to generate more coherent questions using our metrics. Our results show that these interventions can significantly improve VLMs’ accuracy, coherence, and even the explanation efficiency. We lastly show how our multi-faceted metrics for this problem reveal previously impossible insights into VLMs’ reasoning processes, enabling possible targeted improvement in downstream engineering of task guidance systems.
2 Problem Formulation and Dataset
In this section, we define the extended problem of explainable PMD in an approachable manner for VLMs, then we introduce a benchmark dataset we curated for evaluating explainable PMD.
2.1 Defining Explainable PMD
Formally, the inputs for PMD are a short procedural text $T$ and a single video frame $F$, which may or may not visualize the successful completion of the procedure described in $T$. Given these inputs, a system should return a binary mistake detection decision $y$ for whether the procedure has been successfully completed ($y=0$ indicates success, and $y=1$ indicates a mistake). In explainable PMD, it must additionally generate an explanation for the decision $E=(\mathcal{Q},\mathcal{A})$, where $\mathcal{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$ and $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$ are respectively sequences of $n$ yes-no questions and their predicted yes-no answers.
As shown in Figure 1, we instantiate this problem through an iterative self-reflective dialog of questions and answers. This dialog consists of several steps: proposing questions to ask about the procedure at hand (visual question generation or VQG), answering them (visual question answering or VQA), and making a success/mistake classification for whether the procedure has been completed in the given video frame. This structure goes beyond past approaches for PMD using VLMs; while Du et al. (2023) only elicited success classification, Bao et al. (2023) used procedure-agnostic prompts to caption images before making a classification, nonetheless excluding this information from quantitative evaluations.
2.2 Constructing a Dataset for PMD
We follow Du et al. (2023) in recasting Ego4D Grauman et al. (2022), a large-scale egocentric video dataset for everyday activities with dense annotations for various aspects of the videos, into an offline mistake detection format, but expand the diversity of mistake types studied there. Ego4D’s hand and object interactions data subset includes videos of physical actions being performed with various objects. Each video is annotated with narrations describing fine-grained procedures being performed, timestamps for when it begins and ends, and category labels for the verb and noun characterizing the procedure. This makes an ideal testbed for evaluating VLMs’ understanding of real-world actions in video frames, but the data is not formulated for mistake detection. We thus apply several preprocessing steps to the data to create a new Ego4D for Procedural Mistake Detection (Ego4D-PMD) benchmark that includes successful cases and a breadth of mistake types for each annotated procedure, visualized in Figure 2.
Additional details about these preprocessing steps and summary statistics for the Ego4D-PMD dataset are presented in Appendix A. As listed there, to conserve compute, we randomly sample a subset of 10,000, 500, and 2,000 examples respectively from the training, validation, and testing partitions (evenly split between success and mistake cases) for the forthcoming experiments.
Figure 2: Selected examples from Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD). For each matching pair of a video frame and procedural text (in this case, “Fold the cloth with your hands”), we generate a success example, and generate various types of mistake examples by sampling alternate video frames: incomplete execution of the procedure, execution with the wrong verb (e.g., wringing a cloth instead of folding), execution with the wrong noun (e.g., folding a paper instead of a cloth), and execution with both the wrong verb and noun (e.g., opening a notepad instead of folding a cloth). Images slightly cropped for space.
3 Evaluating Coherence in PMD
Next, we describe our application of a fine-tuned NLI model to calculate two coherence metrics for explainable PMD: relevance of questions and informativeness of answers to those questions.
Figure 3: Using BART Lewis et al. (2020) fine-tuned on MNLI Williams et al. (2017) to judge success.
3.1 Using NLI Models to Judge Success
As visualized in Figure 3, LMs fine-tuned for NLI can estimate the sufficiency of visual questions and answers in explaining whether a procedure was successfully completed. To implement this, we need an NLI model $f_{e}$ (which returns a probability that an input premise string entails a hypothesis string), a premise transformation $t_{p}$ (which converts a question $Q$ and answer $A$ into a declarative statement to add to the premise), and a hypothesis prompt template $t_{h}$ (which creates a hypothesis about the success of the procedure in $T$). We can then calculate the NLI model’s probability for the success of procedure $T$ based on the explanation $E=(\mathcal{Q},\mathcal{A})$ as follows:
$p_{e}(T|\mathcal{Q},\mathcal{A})=f_{e}\bigl{(}t_{h}(T)|\{t_{p}(Q_{i},A_{i})|1\leq i\leq n\}\bigr{)}$
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^3] Model weights can be downloaded at https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^4] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^5] While we could also use an LM to rephrase this statement, we found that more complex procedural texts were less likely to be rephrased accurately, and ultimately degraded the correlation with human judgments (see Appendix C).
3.2 Relevance
A coherent decision in PMD should be supported by relevant questions about the state of the environment. [^6] For example, given a procedure “In a bowl, add the cut cherry tomatoes” Peddi et al. (2024), the question “Are there tomatoes in the bowl?” is relevant to the success of the procedure, while the question “Is the bowl blue?” is not. We measure the relevance of a question $Q^{\prime}$ to the success of a procedure $T$, given previous questions $\mathcal{Q}$ and their answers $\mathcal{A}$, as follows:
$\begin{split}\texttt{Rel}(Q^{\prime}|T,\mathcal{Q},\mathcal{A})=&\bigl{|}p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup\text{``{No}''})\\ &-p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup\text{``{Yes}''})\bigr{|}\end{split}$
This definition quantifies how much impact the answer to the proposed question $Q^{\prime}$ can have on the success probability (as estimated by the NLI model). If the success probability is similar for “Yes” and “No” answers, this suggests that $Q^{\prime}$ would not reveal pertinent information (i.e., beyond that in $\mathcal{Q}$ and $\mathcal{A}$) about whether the procedure in $T$ was successfully executed by the user, and thus the relevance would be low. If the success probabilities vary widely depending on the answer, this suggests that $Q^{\prime}$ can reveal important new information to help make the decision.
Example-level relevance.
For evaluation across a dataset, we summarize the relevance of a sequence of questions generated for a particular example of explainable PMD by taking the mean relevance of generated questions with respect to previous questions and answers:
$\frac{1}{n}\sum\limits_{i=1}^{n}\texttt{Rel}\bigl{(}Q_{i}|T,\{Q_{j}:j<i\},\{A_{j}:j<i\}\bigr{)}$
3.3 Informativeness
Since a relevant question does not guarantee an informative answer, [^7] For example, in the procedure “In a bowl, add the cut cherry tomatoes,” “Are there tomatoes in the bowl?” is a relevant question, but a “Yes” answer does not give us enough information to confirm that the procedure is 100% complete (e.g., there could be more tomatoes still outside the bowl). and VLM errors in answering questions could unintentionally introduce conflicting information, it is necessary to evaluate the sufficiency of VLMs’ answers in justifying a PMD decision. To achieve this, we measure the informativeness of a predicted answer $A^{\prime}$ for a question $Q^{\prime}$ to the success of a procedure $T$ (given previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$) as follows:
$\texttt{Inf}(A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A})=1-H\bigl{(}p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup A^{\prime})\bigr{)}$
$H$ is the binary Shannon entropy of the success probability $p_{e}$, calculated by $H(p)=-p\log_{2}p-(1-p)\log_{2}(1-p),p\in[0,1]$. This definition for informativeness quantifies how much information an answer to a question gives us in determining the success of the procedure. As such, if the success probability given this answer $A^{\prime}$ to $Q^{\prime}$ is confident (i.e., very low or high), this indicates that $A^{\prime}$ (along with $\mathcal{Q}$ and $\mathcal{A}$) are sufficient to make a decision, and thus informativeness will be high. On the other hand, a success probability closer to 50% suggests that the information gathered thus far is insufficient, yielding low informativeness. We express informativeness as a percentage (i.e., of 1, the maximum absolute value of bits here).
Reference-adjusted informativeness.
We also wish to account for cases where the information gathered in questions and answers indicates the wrong PMD decision. To do so, we define the NLI model’s PMD belief $y_{e}(T|\mathcal{Q},\mathcal{A})$ as 1 (mistake) if $p_{e}(T|\mathcal{Q},\mathcal{A})<0.5$, else 0 (success). We then define the reference-adjusted informativeness:
$\texttt{Inf}^{*}(A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A},y^{*})=\\ \begin{cases}\texttt{Inf}\bigl{(}A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A}\bigr{)},&y_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup A^{\prime})=y^{*}\\ -\texttt{Inf}\bigl{(}A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A}\bigr{)},&\text{else}\end{cases}$
This makes informativeness negative if the NLI model judges the evidence gathered as indicating the wrong final answer for PMD.
Example-level informativeness.
For evaluation across a dataset, we summarize the (reference-adjusted) informativeness of a set of questions and answers generated for a particular example by taking the maximum value across the self-dialog:
$\max\limits_{1\leq i\leq n}\texttt{Inf${}^{*}$}(A_{i}|Q_{i},T,\{Q_{j}:j<i\},\{A_{j}:j<i\},y^{*})$
4 Experimental Results
In this section, we present a broad set of experiments with VLMs in explainable PMD to validate our proposed metrics. First, we provide details for how VLMs are applied to the problem. We then introduce two inference-time interventions to encourage VLMs to select more coherent questions from generated candidates, as well as an approach to encourage them to generate more coherent candidates through preference optimization from our proposed coherence metrics. After evaluating and comparing these approaches, we show that our proposed metrics visualize and characterize common behaviors of VLMs, enabling a panoptic understanding of their capabilities in PMD.
4.1 Applying VLMs to Explainable PMD
Our experiments evaluate explainable PMD in InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024). [^8] See https://huggingface.co/Salesforce/instructblip-vicuna-7b, https://huggingface.co/liuhaotian/llava-v1.5-7b, and https://huggingface.co/meta-llama/Llama-3.2-11B-Vision. We intentionally target smaller, open-source VLMs that are feasible for online task guidance. The models apply different architectures, training datasets, and/or training strategies to integrate vision into the model. While LLaVA and InstructBLIP were not trained on Ego4D data, Dubey et al. (2024) does not reveal whether Llama 3 was. Details and prompt templates can be found in Appendix D, while key implementation details about the iterative self-dialog are provided below.
VQG.
We first prompt the VLM to generate a series of questions given the procedural text and the dialog history with previous questions and answers (enabling deductive reasoning). To encourage logical but diverse questions, we apply greedy beam search decoding to generate 4 candidates, from which we select the most likely candidate that has not already been generated.
VQA.
After a question is generated, it is answered by the VLM given only the question and video frame. If the probability of the chosen answer (i.e., “Yes” or “No”) exceeds an answer sureness threshold of 60%, we append it to the dialog history, otherwise we append “Unsure.” These unsure answers are excluded from the sets of questions $\mathcal{Q}$ and answers $\mathcal{A}$ in the calculation of example-level relevance and informativeness (defined in Section 3). Effectively, they are assigned zero informativeness and excluded from previous questions and answers in metric calculations.
Success classification.
After each iteration of VQG and VQA, we prompt the VLM to judge whether the procedure has been successfully executed based on the video frame and entire dialog history. The VLM’s decision is made using a mistake confidence threshold $\tau$ (tuned on the validation data for each approach) on its mistake likelihood. The prompt and answer for this step are excluded from the dialog history in future iterations.
Stopping criteria.
To prevent over-generating noisy information, which can degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before $n$ questions have been generated) if the likelihood of success either stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). The hyperparameters $n$, $\delta$, and $\epsilon$ are tuned based on the validation data for each presented approach.
4.2 Coherent Question Selection
While a typical beam search would rank candidate questions by their likelihood, an alternative approach is to re-rank candidates using the reference-free coherence metrics introduced in Section 3. This could encourage the selection of questions that are most likely to bring in new, salient, and helpful information. Furthermore, as prior work has shown how the in-context learning capability of LMs enables them to accurately predict physical states from just a few in-context demonstrations Zhang et al. (2023), we will examine how supplementing candidate questions generated from the dialog history with additional questions generated through in-context learning approach impacts VLM performance.
Next, we introduce these two approaches we use to augment the candidate question pool for more coherent candidates: coherence-based re-ranking and candidate generation through in-context learning. We then compare their performance on Ego4D-PMD with that of vanilla VLMs.
4.2.1 Coherence-Based Question Selection
We implement a coherence-based candidate question re-ranking approach as follows. Given a set of question candidates $\hat{\mathcal{Q}}$ for procedural text $T$ along with previous confidence-filtered questions $\mathcal{Q}$ and answers $\mathcal{A}$, we can select the best question $Q^{*}$ by maximizing the product of relevance and potential informativeness across all $Q\in\hat{Q}$:
$\texttt{Rel}(Q|T,\mathcal{Q},\mathcal{A})\max\limits_{A\in\{\text{{Yes}},\text{{No}}\}}\texttt{Inf}(A|Q,T,\mathcal{Q},\mathcal{A})$
This ranking prioritizes well-rounded questions that can yield impactful information for success classification which leads to the most confidence. $Q^{*}$ is then concatenated to the dialog history and answered by the VLM as described earlier.
InstructBLIP
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 65.5 19.0 25.1 2.99 0.266
L 66.8 16.5 31.1 3.64 0.323
C 64.4 27.0 29.8 3.54 0.290
C 67.4 39.6 32.1 2.56 0.306
LLaVA
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.6 41.9 24.0 2.77 0.427
L 65.0 36.2 28.1 3.75 0.438
C 65.0 65.3 33.3 3.32 0.566
C 68.1 74.2 46.0 3.43 0.674
Llama 3
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.3 17.5 27.0 4.82 0.284
L 57.7 19.6 15.8 1.96 0.194
C 63.7 27.1 32.0 5.81 0.325
C 62.7 52.4 45.1 3.72 0.418
Table 1: Ego4D-PMD test set results for likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
LLaVA + DPO
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 63.5 94.5 29.2 1.72 0.822
L 64.4 67.3 41.0 2.98 0.723
C 61.5 95.9 30.8 1.87 0.782
C 61.9 97.0 30.4 1.75 0.778
Table 2: Ego4D-PMD test set results for LLaVA with coherence-based fine-tuning through DPO.
4.2.2 In-Context Learning Augmentation
Applying in-context learning in explainable PMD is not straightforward, as it would require reasoning over multiple images and possibly long, variable-length dialogs. Instead, we use in-context learning to improve the text-based VQG step by providing examples of human-written questions. We achieve this by annotating 20 procedures from the Ego4D training data with 3 reasonable questions one could ask about a given procedure to judge its success. [^9] These human-written questions achieve a mean 53.9% relevance and 83.2% maximum informativeness (i.e., for a yes or no answer). Procedures and questions in Appendix E.1. We prompt the VLM with these example procedures and questions, the current procedure at hand, and the previous 2 questions proposed by the VLM (as available) to incorporate information the VLM already collected. We then generate 4 additional candidate questions using the same constraints as described earlier. To minimize the impact of ordering, in-context examples are randomly shuffled in every prompt; a random seed is used to ensure examples are equivalently shuffled in different experiments.
4.2.3 Comparison of Results
We compare these approaches to vanilla VLMs based on mistake detection accuracy, mean example relevance and reference-adjusted informativeness of generated explanations (as defined in Section 3). We add two metrics for the efficiency of the self-dialog: the average number of iterations that occurred before stopping (i.e., the length of the dialog), and the average information gain in the success likelihood across all iterations in bits (i.e., how much information the VLM got from it).
Experimental results are presented in Table 2.101010Selected hyperparameters and validation set results for this table are listed in Appendix E.2. We find that coherence-based re-ranking and in-context learning [^11] To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources in Appendix E.3. both sharply improve the relevance and informativeness in all models, reaching a respective 74.2% and 46.0% in LLaVA. This suggests that questions that VLMs judge to be most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 68.1%, demonstrating that coherent explanations are valuable to accurate PMD. Lastly, we find that the information gain under our proposed question selection interventions is consistently higher, reaching a maximum of 0.674 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent explanations.
4.3 Coherent Question Generation
While the above approaches boost the coherence of VLM outputs for PMD, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. As such, we lastly explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
Specifically, we apply direct preference optimization (DPO; Rafailov et al., 2023) with quantized low-rank adaptation (QLoRA; Hu et al., 2022; Dettmers et al., 2024) to fine-tune a VQG adapter for LLaVA (the best performing model). We generate training data for preference optimization by first running inference over the sampled Ego4D-PMD training data with coherence-based candidate question re-ranking and additional candidates generated through in-context learning. [^12] Here, we use a default set of inference hyperparameters $n=10$, $\delta=0.1$, and $\epsilon=0.05$. An ablation without in-context learning is presented in Appendix E.4. We then fine-tune VLMs on pairs of chosen and rejected candidate questions for each self-dialog turn based on their coherence-based ranking. [^13] Additional details, selected hyperparameters, and validation set results are listed in Appendix E.5. While the top ranked question is always used as the chosen question, the rejected question is randomly sampled from the bottom half of ranked questions. During explainable PMD evaluation, the trained adapter is applied to the VLM during VQG steps, and disabled for other steps.
4.3.1 Comparison of Results
Table 2 displays the results of fine-tuning LLaVA for VQG, including the previous inference-time interventions to compare their relative impact on fine-tuned models. First, we observe that the relevance drastically increases from a maximum of 74.2% to 97.0%. Remarkably, this shows that VLMs are capable of learning to ask more coherent questions for explainable PMD, a task based on properties of the physical world.
However, we find that accuracy and informativeness drop slightly from the best results in Table 2, suggesting that asking more coherent questions is not enough to improve performance globally. We suspect that this is due to an implicit difficulty trade-off of asking highly relevant questions, which should strongly indicate success if answered one way, otherwise a mistake. For elaborate procedures, highly relevant questions may cover multiple aspects or states of the scene. As such, answering these questions may be more difficult for VLMs, which struggle with complex questions. For example, while the generated question “Is the soil around the tomato seedling with the gardening trowel in your hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering such a question requires understanding spatial relations between several objects (i.e., soil, tomato seedling, trowel, and hand). Future work may explore methods to encourage VLMs to generate simpler questions.
Meanwhile, we find that fine-tuning reduces the average number of iterations taken to come to a conclusion by about 1, reaching a minimum of 1.72. Further, the average information gain reaches a new peak of 0.822. This suggests that coherence-based fine-tuning can empower VLMs to make more confident decisions faster. We lastly observe that the improvement in all metrics from coherence-based ranking and in-context learning is generally smaller in fine-tuned LLaVA than the base model (presented in Table 2). This suggests that our coherence metrics can improve the effectiveness of base VLMs for explainable PMD.  [^14] For a reference point with proprietary LMs, see Appendix E.6 for an evaluation of GPT-4o OpenAI et al. (2024).
4.4 Visualizing PMD Performance
An additional benefit of our automated coherence metrics is the ability to audit the global and local reasoning behaviors of VLMs under various strategies. In Figure 4, we visualize the distribution of decision error, relevance, and informativeness of four representative approaches to apply LLaVA. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based fine-tuning with DPO.
For each example, decision error is calculated by how far the VLM’s success likelihood was from being 100% confident in the ground truth PMD decision. Relevance and informativeness use the example-level forms defined in Section 3. The colors of points indicate combinations of decision error, relevance, and informativeness in VLM outputs, highlighting common outcomes; we describe several examples in Appendix E.7. In comparing the plots, we see a dramatic shift as each successive intervention is introduced. Specifically, we see that red and black/indigo points are virtually eradicated, indicating fewer incoherent explanations and complete failures.
Vanilla LLaVA
+ Coherence Ranking
+ ICL
+ DPO
Figure 4: Visualization of decision error, relevance, and reference-adjusted informativeness for configurations of LLaVA on Ego4D-PMD testing examples. Each data point’s color indicates its position along each axis.
Ultimately, when choosing an approach to use for explainable PMD, one should weigh accuracy, coherence, and explanatory dialog efficiency as appropriate for the setting. For example, online applications may benefit from shorter (i.e., faster) dialog, even at a cost of accuracy. Meanwhile, high-risk applications may prioritize high confidence (e.g., information gain) over other factors. Additional insights toward the fine-grained strengths and weaknesses of various approaches may be gained from analyzing these results by mistake type, or verb and noun categories.
5 Related Work
In Section 1, we contextualized this work with prior work in PMD. In this section, we additionally discuss two lines of work: multi-step reasoning in LMs and leveraging NLI models in NLP.
5.1 Multi-Step Reasoning in LMs
Foundational LMs have exhibited impressive capabilities in multi-step reasoning from prompting-based methods Wei et al. (2022); Kojima et al. (2022). Later work strengthened LMs’ reasoning with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024a); Qi et al. (2024), and fine-tuning on data generated by stronger LMs Wang et al. (2024); Gou et al. (2024). Related to task guidance, some work has investigated LMs’ ability to reason about dependencies between steps of physical procedures Bellos et al. (2024); Lal et al. (2024). In light of new challenges in visually-dependent reasoning Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024); Zhang et al. (2024b), recent work has strengthened VLMs by utilizing text-based LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024). In this work, we similarly enabled VLMs to coordinate their visual reasoning through a self-dialog in the novel problem of explainable PMD, and we additionally proposed new coherence metrics to evaluate and improve their reasoning.
5.2 Leveraging NLI in Other NLP Tasks
NLI requires determining whether a premise text entails a hypothesis text, and has long been studied in the NLP community Dagan et al. (2005). Once thought to be a grand challenge for commonsense reasoning, many human-annotated resources have been compiled for NLI, and thus significant progress has occurred in training specialized LMs for NLI Storks et al. (2019). Prior work has leveraged this progress by using such NLI models to improve the competence, confidence, and coherence of LMs in other tasks, such as dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), visual question answering Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). Our work extends this effort to a novel problem of explainable PMD, adopting two NLI-based coherence metrics.
6 Conclusion
In this work, we investigated the capabilities of foundational VLMs in a challenging new problem of explainable procedural mistake detection (PMD), where visual questions and answers must be generated to drive system decisions. To evaluate explanations, we defined two coherence metrics leveraging a fine-tuned NLI model. We then applied several interventions in VLMs to encourage coherent question selection (e.g., coherence-based question re-ranking) and generation (e.g., preference optimization from pairs of coherent and incoherent questions). Our results showed that vanilla VLMs do not generate coherent explanations off-the-shelf, but these interventions vastly improve their coherence, with the former also improving their overall PMD accuracy, and the latter improving the efficiency of generating and extracting information from explanations. Further, we showed that patterns in accuracy and coherence metrics illuminate common failures, e.g., visual hallucination. This work lays a foundation for future work harnessing VLMs for task guidance and other multimodal reasoning applications.
7 Limitations
Latency of self-dialog.
One limitation of our explainable PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings. This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.
Inherent limitations of single video frames.
Next, single video frames are limited in representing actions, which involve movement and state change. Our decision to focus on individual frames stemmed from preliminary experiments we performed with existing open-source VLMs for video understanding Lin et al. (2024); Li et al. (2023a), which are still in early stages. There, we found that they often confused information from frames in different regions of the video, preventing them from judging the final states of objects and reconciling this information with success of procedures, and resulting in poor performance. As such, our choice to focus on single images simplified the problem for current VLMs, enabling our experiments to begin building a meaningful understanding of their capabilities in PMD. To minimize the dependence on multiple frames in detecting mistakes, we applied several careful preprocessing steps to Ego4D-PMD (as discussed thoroughly in Section 2.2 and Appendix A). It is also worth noting that by not incorporating modalities beyond text and images, e.g., audio, the VLMs we studied are inherently limited in their capturing of physical information Yu et al. (2022); Zong et al. (2024). As the ability to reason over sequences of frames and other modalities evolves in state-of-the-art VLMs, future work can revisit this formulation and explore new approaches for them to reason over dynamic scenes.
Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003. We would like to thank Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, as well as the entire MSRP team for their helpful discussions, feedback, and insights.
Appendix A Ego4D-PMD Data Curation Details
Various benchmark datasets have been created for PMD from egocentric video, each of which includes video and procedural text along with various other modalities, as well as detailed information about mistakes Du et al. (2023); Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024). While these datasets are useful resources for research in task guidance, they are video-based and cover limited domains. Further, most of them include dialog interaction between a user and instructor agent which often causes mistakes to be corrected before or while they happen. They include mistakes around temperature, timing, small measurements, and other physical properties of the environment that are difficult for open-source VLMs, which are mostly optimized for representing single images, to perceive. While some of these issues could be overcome through a two-way dialog between the agent and a user, this makes it harder to isolate mistakes occurring in the videos and dive deep into the reasoning behind detecting them.
To alleviate these challenges and focus our inquiry, we follow Du et al. (2023) in recasting Ego4D, a procedural video dataset with a breadth of narrated everyday actions, into a mistake detection format. The collection of the Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) data consisted of several preprocessing steps, outlined below.
Generating success examples.
As discussed above, Ego4D’s hand object interaction data is annotated in units of egocentric video clips of individual actions being performed by humans. We can form an example of a successful execution of the procedure by pairing each video clip with its annotated natural language narration of the procedure. Since most VLMs are not optimized to reason over multiple frames and videos, and those that are are still in very early stages, we sample exactly one frame from each video clip. Specifically, as each clip is carefully annotated with a postcondition time for the action, i.e., the time that the action has been completed, we simply sample the video frame at this annotated time and pair it with the text narration.
Generating mistake examples for incomplete procedures.
One natural type of mistake a user could make is not finishing a procedure. In addition to postcondition times, each video clip is annotated with a precondition time. Following a similar approach in Du et al. (2023), we can generate a mistake example by sampling a frame at the precondition time and pairing it with the video clip’s narration text. We expect that by doing this, the sampled frame will show the procedure at an incomplete state, and contain most of the same objects as the success example for the same clip. This poses a difficult challenge of identifying the key physical properties of the scene that would indicate completion.
Generating mistake examples for mismatched verbs and nouns.
Mistakes also happen when a user applies the wrong type of action to an object, causing an unexpected state, as well as when a user uses the wrong object or ingredient in a procedure. Following this intuition, we generate additional easier mistake examples from each clip by matching each clip with other clips that have a mismatched verb, noun, or both. While each clip is annotated with verb and noun categories, these categories are coarse-grained, making it impossible to guarantee that two clips with the same verb or noun label actually involve the same verb or noun, thus preventing sampling clips that share the same verb or noun. Instead, we apply the AllenNLP [^15] https://allenai.org/allennlp semantic role labeler to each narration text to identify the key participants in each procedure. For each clip, we then attempt to sample the postcondition frames from three mismatched clips: one with a mismatched verb (but matching nouns), one with a mismatched noun (but matching verb), and one with a mismatched verb and noun. We then pair these frames with the source clip’s narration text, creating mistake examples with varying levels of overlap with the source clip. While it is not always possible to find every such alternative clip for each clip in Ego4D, we can usually find at least one of them.
Transforming narrations into instructions.
The narration texts annotated in Ego4D are declarative statements about the actions being performed in each clip. This is not an accurate depiction of typical interactive task guidance and PMD settings, which usually revolve around instructional texts like recipes or guidebooks. As such, we convert each narration, e.g., “Someone washes the lettuce,” into imperative form, e.g., “Wash the lettuce,” using spaCy. [^16] https://spacy.io/ Further, some narrations describe procedures that are not suited for comparing physical state changes in text and images, such as social interactions, interactions with animals, interactions with electronic devices, and movements that are impossible to precisely characterize from the narration text (e.g., in “Move plate”). We use the verb and noun category annotations on each clip to filter out such cases.
Ensuring data quality.
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed over and over (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for explainable PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a handful of videos in Ego4D that we notice to be corrupted or significantly distorted.
Various statistics of the resulting full Ego4D-PMD dataset (and sub-samples used in the paper) are presented in Tables 3 and 4. It is worth noting that this collected dataset does not include ground truth explanations for mistake detection labels (rather we proposed automated reference-free metrics) for two important reasons. First, there may be multiple valid ways to detect a mistake through asking and answering visual questions, each of which could consist of different questions and/or different numbers of questions. [^17] For example, in trying to determine the success of the procedure “In a bowl, add the cut cherry tomatoes” from Peddi et al. (2024), we could reasonably ask one question “Are all the cherry tomatoes in the bowl?” or two questions “Are there cherry tomatoes in the bowl?” and “Are there any cherry tomatoes outside of the bowl?” In our opinion, existing metrics for text generation are not as well suited for this extremely challenging evaluation as fine-tuned NLI models, which are optimized to judge the logical consistency between texts. Second, in a real-world setting like PMD for task guidance, we believe that automated metrics are better suited for understanding and improving a deployed system than an offline benchmark of ground truth explanations.
Type Train (Sample) Val. (Sample) Test (Sample) Total
Success 42.0k 5.00k 13.1k 250 18.1k 1.00k 73.1k
Mistake 99.4k 5.00k 25.4k 250 34,182 1.00k 159k
(Incomplete) 15.1k 755 4.91k 51 6.55k 194 26.5k
(Wrong V) 11.8k 604 2.69k 31 3.75k 108 18.2k
(Wrong N) 36.4k 1.85k 8.91k 87 11.8k 344 57.2k
(Wrong V&N) 36.1k 1.79k 8.91k 81 12,047 354 57.1k
Table 3: Distribution of example types in each partition of our curated Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) dataset.
Type Train (Sample) Val. (Sample) Test (Sample) All
Verbs 83 80 77 55 78 71 83
Nouns 440 372 365 151 390 257 487
V-N Pairs 3,976 2,050 2,185 326 2,658 833 5,363
Table 4: Distribution of unique verb, noun, and verb-noun pair categories in each partition of our curated Ego4D-PMD dataset. Verb and noun categories are annotated for each narration in the Ego4D dataset.
Appendix B Rephrasing Questions and Answers for Coherence Evaluation
As discussed in Section 3.1, we use a fine-tuned NLI model to judge the success of procedures given questions and answers. In order to convert questions and answers into declarative statements to pass into the NLI model, we follow Srinivasan et al. (2024) in prompting a VLM with the following 10 in-context demonstrations of rephrasing before prompting it to rephrase a question and answer for the task at hand:
  1. 1.
    Question: Is there a bowl on the table?
    Answer: Yes
    Statement: There is a bowl on the table.
  2. 2.
    Question: Are the eggs cracked?
    Answer: No
    Statement: The eggs are not cracked.
  3. 3.
    Question: Does the cardboard box look open?
    Answer: Yes
    Statement: The cardboard box looks open.
  4. 4.
    Question: Are there any leaves outside of the basket?
    Answer: No
    Statement: There are not any leaves outside of the basket.
  5. 5.
    Question: Is the orange peeled?
    Answer: Yes
    Statement: The orange is peeled.
  6. 6.
    Question: Is the mug empty?
    Answer: No
    Statement: The mug is not empty.
  7. 7.
    Question: Are there hedge trimmers in the image?
    Answer: Yes
    Statement: There are hedge trimmers in the image.
  8. 8.
    Question: Has the light switch been turned on?
    Answer: No
    Statement: The light switch has not been turned on.
  9. 9.
    Question: Does the table have any cups on it?
    Answer: Yes
    Statement: The table has cups on it.
  10. 10.
    Question: Is the cabinet closed?
    Answer: No
    Statement: The cabinet is not closed.
Appendix C Coherence Metrics Human Study Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3. To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of evaluated approaches:
  • Likelihood-based question ranking
  • Coherence-based question ranking augmented with question candidates from in-context learning
C.1 Annotation Instructions
For both relevance and informativeness annotation, we provided the following background for the task:
Imagine you just had eye surgery, and are currently unable to see. You’re performing a task you’re familiar with, but need help to determine whether you successfully completed it. You video call a friend (who is unfamiliar with the task) and show them what you’re working on. You then ask them some yes/no questions to figure out whether you successfully completed the task.
Relevance annotation instructions and example.
Annotators are provided the following instructions for annotating relevance:
For each annotation task, you will be given the following information:
  • A sentence describing the procedure you’re trying to perform.
  • An optional list of previous questions you already asked, and their answers.
  • A potential next question you could ask your friend.
You must rate how relevant the potential next question is. By relevant, we mean: given the previous questions and answers, how helpful could an answer to this question be in determining whether you successfully completed the task?
We recruit 5 annotators, each of which annotates 10 samples. One sample is listed below:
Sentence: Drop the bowls on the table with your hand
Previous questions and answers:
  1. 1.
    Are the bowls on the table? (Answer: Yes)
  2. 2.
    Is the person holding the bowls in their hand? (Answer: No)
Potential next question: Is the person about to drop the bowls on the table?
Your rating:
  • 1 (very irrelevant)
  • 2 (slightly irrelevant)
  • 3 (neutral; may or may not be relevant)
  • 4 (slightly relevant)
  • 5 (very relevant)
  • Instructions Unclear
Informativeness annotation instructions and example.
Annotators are provided the following instructions for annotating informativeness:
For each annotation task, you will be given the following information:
  • A sentence describing the procedure you’re trying to perform.
  • A list of questions you asked your friend, and their answers.
You must rate how informative the questions and answers are. By informative, we mean: based on all the information you have, how sure are you about whether you succeeded?
We recruit 5 annotators, each of which annotates 10 samples. One sample is listed below:
Sentence: Clean the bowl
Previous questions and answers: None
Last question: Is there a bowl in the image? Last answer: Yes
Your rating:
  • 1 (very uninformative/unsure)
  • 2 (slightly uninformative/unsure)
  • 3 (neutral; may or may not be relevant)
  • 4 (slightly informative/sure)
  • 5 (very informative/sure)
  • Instructions Unclear
C.2 Human Study Results
For each metric, we recruited 5 annotators (all English speakers with with conferred or in-progress undergraduate degrees) to rate the relevance and informativeness of machine-generated questions and answers.
Comparison of automated and human informativeness judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., given the previous questions and answers, how helpful could an answer to this question be in determining whether the task was successfully completed) on a scale from 1-5 (least to most relevant). Between the resulting 50 annotations and corresponding automated metrics, we found a moderate Spearman correlation Spearman (1904) of $\rho=0.55$ ($p=0.000029$). This suggests that this automated measure of relevance is indeed positively correlated with human judgments of relevance.
Comparison of automated and human informativeness judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.32$ ($p=0.024$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $Q^{\prime}$, $\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.2.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 4.4, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.
Appendix D VLM Self-Dialog Details
In this appendix, we include prompt templates and other supplementary details for the self-dialog we conditioned VLMs to generate, including visual question generation, visual question answering, and stopping criteria. To conserve GPU memory, VLM weights are 4-bit quantized at inference time.
D.1 Visual Question Generation
When prompting vision-and-language models (VLMs) to generate questions, we use the following prompt template:
[…]
Q:
Question generation is not conditioned on video frames, as we found significant performance degradation when VLMs were conditioned on the video frame while generating questions, often leading to completely nonsensical questions, e.g., “Is is is is is is?” To ensure the VLM generates yes-no questions, we constrain generation during decoding to enforce that each generated text begins with a word that can signal a yes-no question, [^18] Specifically, questions must begin with is, does, or has, along with all plural and past tense forms of these verbs. does not include the word or, and ends with a question mark. To avoid vague and high-level questions about the status of the procedure rather than low-level physical states, we prevent VLMs from repeating some words present in their input prompts: “successful,” “successfully,” “executed,” and “procedure.” To encourage logical questions while ensuring variety, we apply greedy beam search decoding with $k=8$ beams, returning the top 4 candidate questions to be ranked through LM likelihood or coherence metrics. [^19] Due to generation constraints, it is often the case that the VLM does not successfully generate all 8 candidates. In rare cases, these constraints prevent VLMs from generating any candidate questions; when this happens, we repeat the generation without any constraints (even though this may result in outputs that are not yes-no questions). Out of the candidates, we remove any that are exactly the same as previously generated questions, then select the most likely candidate based on the model’s log-likelihoods.
D.2 Visual Question Answering
Once a question is generated, we prompt the VLM with it and the video frame along with “A:” to elicit an answer. To produce the answer, we apply softmax over the output logits from the forward pass of the VLM for the Yes and No tokens. It is important to note that we exclude the dialog history from the context during VQA, as we again observed significant performance degradation when VLMs answered visual questions in the context of a longer dialog. This was especially prominent when several similar questions were generated and answered in a dialog, which often caused the VLM to creep from being initially unsure about the answer to being confidently wrong.
D.3 Success Classification
Questions and answers are generated iteratively until the stopping criteria described in Section 4.1 are met. To prompt the VLM to judge the success of a procedure after a question is answered, we use the following prompt:
[…]
Here, the logits of the Yes and No tokens are similarly used to produce a probability distribution over a success or mistake decision. The decision is determined by a mistake confidence threshold $\tau$, which is selected to maximize overall accuracy from a comprehensive set of 99 candidates $\tau\in\{0.01,0.02,\dots,0.98,0.99,1.0\}$.
D.4 Stopping Criteria Details
These hyperparameters $n$, $\delta$, and $\epsilon$ for the self-dialog stopping crtieria are selected from a grid search on the validation data over combinations of $n\in\{2,4,6,8,10\}$, $\delta\in\{0.05,0.1,0.2,0.4\}$, and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a summary metric which is the product of example-level informativeness (reference-adjusted) and relevance if and only if the VLM makes an accurate mistake detection prediction and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described above. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.2, while values for the results in Table 2 are listed in Appendix E.5.
Appendix E Experiment Details and Supplementary Results
In this appendix, we provide assorted details (e.g., hyperparameters) and supplementary results that were omitted from the main body of the paper.
E.1 Example Questions for In-Context Learning in Question Generation
As discussed in Section 4.2.2, we condition VLMs with sets of human-written questions for 20 procedures from the Ego4D for Procedural Mistake Detection (Ego4D-PMD) dataset. The annotated procedures (underlined) and questions (italicized) are listed below:
  1. 1.
    Soak the sponge in a soapy water with your hands
    1. (a)
      Is there a sponge?
    2. (b)
      Is the sponge in water?
    3. (c)
      Is the water soapy?
  2. 2.
    Open the bottle
    1. (a)
      Is there a bottle in the image?
    2. (b)
      Is the bottle open?
    3. (c)
      Does the bottle have a lid on it?
  3. 3.
    Take the baking tray away from the table
    1. (a)
      Can you see a baking tray?
    2. (b)
      Is the baking tray on the table?
    3. (c)
      Is the baking tray picked up by someone?
  4. 4.
    Turn on a torch light
    1. (a)
      Is there a torch light in the photo?
    2. (b)
      Is the torch light powered on?
    3. (c)
      Is the torch light lit up?
  5. 5.
    Fold the right edge of the wrapper
    1. (a)
      Is there a wrapper in the image?
    2. (b)
      Is the wrapper completely flat?
    3. (c)
      Is the right edge of the wrapper folded?
  6. 6.
    Pour the water into the blue container
    1. (a)
      Do you see a blue container anywhere?
    2. (b)
      Is there water in the blue container?
    3. (c)
      Is the blue container empty?
  7. 7.
    Paint the patio with the paint brush
    1. (a)
      Is this a photo of a patio?
    2. (b)
      Is the patio painted?
    3. (c)
      Is someone holding a paint brush?
  8. 8.
    Spread the black peas on the salad with the spoon in your hand
    1. (a)
      Is there a salad?
    2. (b)
      Are there black peas on the salad?
    3. (c)
      Is there a spoon in someone’s hand?
  9. 9.
    Scoop paint from the pallet on the table with the paint brush
    1. (a)
      Do you see a paint brush and a paint palette?
    2. (b)
      Is there paint on the paint brush?
    3. (c)
      Is the paint brush in someone’s hand?
  10. 10.
    Wash the car with a sponge in your hand
    1. (a)
      Do you see a car?
    2. (b)
      Is the car clean?
    3. (c)
      Is the sponge being held?
  11. 11.
    Pick the scrubber from the sink
    1. (a)
      Do you see a scrubber somewhere?
    2. (b)
      Is the scrubber in the sink?
    3. (c)
      Is the scrubber in someone’s hand?
  12. 12.
    Peel the onion
    1. (a)
      Is there an onion in the image?
    2. (b)
      Is the onion’s skin removed?
    3. (c)
      Is the onion peeled?
  13. 13.
    Put the dirt in the dust bin
    1. (a)
      Is there a dust bin?
    2. (b)
      Is there dirt in the dust bin?
    3. (c)
      Is there any dirt outside of the dust bin?
  14. 14.
    Cut dough into two
    1. (a)
      Do you see any dough?
    2. (b)
      Is the dough in two pieces?
    3. (c)
      Is the dough whole?
  15. 15.
    Break the walnut with the nutcracker in your hand
    1. (a)
      Do you see a walnut?
    2. (b)
      Is the walnut cracked?
    3. (c)
      Is there a nut cracker in someone’s hand?
  16. 16.
    Turn off the tap
    1. (a)
      Is there a tap in the photo?
    2. (b)
      Is the water running?
    3. (c)
      Is the faucet switched off?
  17. 17.
    Heat the edge of the bag with the lighter
    1. (a)
      Do you see a bag and a lighter?
    2. (b)
      Is there a flame coming from the lighter?
    3. (c)
      Is the lighter near the bag?
  18. 18.
    Close the fridge
    1. (a)
      Is there a fridge?
    2. (b)
      Is the fridge open?
    3. (c)
      Can you see inside the fridge?
  19. 19.
    Chop green beans with a knife on the chopping board
    1. (a)
      Do you see green beans on a cutting board?
    2. (b)
      Are the green beans sliced?
    3. (c)
      Is someone using a knife?
  20. 20.
    Drop the brush in your hand on the oven
    1. (a)
      Is there a brush in the scene?
    2. (b)
      Is there an oven?
    3. (c)
      Is the brush on the oven?
E.2 Question Selection Inference Hyperparameters and Validation Results
In Table 5, we list the inference hyperparameters for the question selection results presented in Table 2: number of iterations $n$, early stopping $\delta$ and $\epsilon$, and mistake confidence threshold $\tau$. In Figure 5, we use detection error tradeoff (DET) curves to visualize the range of accuracy achieved with all candidate mistake confidence thresholds $\tau$ for the approaches compared in Table 2. The full validation set results with selected hyperparameters are listed in Table 6.
InstructBLIP
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 4 0.05 0.2 0.28
L 4 0.05 0.1 0.27
C 10 0.05 0.2 0.42
C 6 0.4 0.2 0.35
LLaVA
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 4 0.05 0.1 0.73
L 6 0.05 0.05 0.73
C 6 0.05 0.05 0.42
C 10 0.05 0.05 0.19
Llama 3
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 10 0.1 0.025 0.31
L 2 0.05 0.1 0.60
C 8 0.05 0.05 0.45
C 6 0.2 0.025 0.39
Table 5: Selected inference hyperparameters for the results presented in Table 2.
InstructBLIP
LLaVA
Llama 3
Figure 5: Mistake detection error tradeoff (DET) curves for VLMs applied to the Ego4D-PMD validation set with likelihood- and coherence-based candidate question selection approaches, with optional supplementary candidates generated through in-context learning (ICL).
InstructBLIP
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 61.6 19.4 26.5 2.93 0.267
L 62.8 16.9 30.9 3.59 0.313
C 64.0 26.7 29.6 3.51 0.295
C 65.2 39.9 24.7 2.51 0.327
LLaVA
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.0 43.1 24.9 2.79 0.424
L 62.0 39.4 28.9 3.68 0.434
C 63.2 66.7 31.9 3.32 0.562
C 64.2 76.6 41.6 3.50 0.663
Llama 3
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 63.6 18.0 28.6 4.93 0.266
L 63.0 19.5 16.0 1.95 0.189
C 65.6 25.5 32.9 5.80 0.331
C 64.4 52.6 48.8 3.62 0.435
Table 6: Ego4D-PMD validation set results for likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.3 Analysis of Question Sources in In-Context Learning
To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources on the validation data in Figure 6. As expected, candidates generated with in-context learning are only rarely selected in the likelihood-based ranking, amounting to about 26.3% of VQG iterations for InstructBLIP, 9.0% of VQG iterations for LLaVA, and 4.9% of VQG iterations for Llama 3. On the other hand, they are selected more frequently in the coherence-based ranking, amounting to about 35.3% of VQG iterations for InstructBLIP, 38.0% of VQG iterations for LLaVA, and 36.5% of VQG iterations for Llama 3. Interestingly, in-context learning candidates are more dominant in earlier iterations, while candidates generated based on the dialog context are relatively more common in later iterations. This may suggest that after selecting a few questions from in-context learning in earlier iterations, the VLM is able to utilize them to generate better questions from the dialog context in later iterations. Alternatively, this could suggest that candidates from in-context learning have limited variety, and thus are less likely to be selected in later turns to avoid redundant questions or information.
InstructBLIP
LLaVA
Llama 3
Figure 6: Histograms of VLMs’ selected question sources, either explanatory dialog context or in-context learning (ICL) examples, by visual question generation (VQG) iteration for likelihood-based question selection (top) and coherence-based question selection (bottom). For a controlled comparison in this visualization, we fix the stopping criteria $n=10$, $\delta=0.1$, and $\epsilon=0.05$.
E.4 DPO Training Data Composition Ablation
In Table 7, we include an additional result for DPO where generated training data does not include candidate questions generated with in-context learning. As shown, compared to the results in Table 2, all metrics reach slightly slower maxima in this ablated configuration.
LLaVA + DPO (without ICL)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 61.9 88.0 30.0 1.82 0.731
L 64.2 65.6 30.8 2.18 0.617
C 61.8 94.8 33.4 1.83 0.761
C 62.3 94.9 34.5 1.88 0.764
Table 7: Ego4D-PMD test set results for DPO-trained VLMs without applying in-context learning in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.5 Question Generation Fine-Tuning and Inference Details
[…]
LLaVA + DPO
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.1 8 0.1 0.05 0.50
L 6 0.05 0.025 0.38
C 6 0.1 0.05 0.04
C 4 0.05 0.05 0.08
LLaVA + DPO (without ICL)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.5 4 0.4 0.05 0.67
L 4 0.4 0.05 0.67
C 4 0.05 0.1 0.77
C 4 0.05 0.05 0.77
Table 8: Selected training and inference hyperparameters for the results presented in Tables 2 and 7.
LLaVA + DPO
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 64.6 94.2 30.8 1.80 0.819
L 66.4 75.5 38.4 2.65 0.759
C 61.0 95.9 27.1 1.89 0.775
C 60.8 96.7 27.1 1.79 0.768
LLaVA + DPO (without ICL)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.6 87.7 29.6 1.81 0.720
L 62.8 72.0 30.6 2.07 0.651
C 65.2 94.4 34.6 1.82 0.748
C 64.8 94.5 34.4 1.88 0.747
Table 9: Ego4D-PMD validation set results for DPO-trained VLMs both with and without applying in-context learning in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.6 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024) [^20] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for explainable PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
To enable the GPT-4o evaluation, we make a few small changes to the VLM self-dialog implementation used in the main experiments. Since GPT can not be forced to generate yes-no questions with specific generation constraints, as done in our open-source implementation, we slightly modify the VQG prompt to encourage appropriate question generation. We add the following to the end of our original VQG prompt, with bold indicating the new text:
This is…from the photo. Generate an appropriate yes/no question.
Q:
To produce a probability distribution over the Yes and No tokens, used in both the VQA and success classification parts of the VLM self-dialog, we use the log probabilities provided by the API as logits are not directly available. We then normalize the probabilities to get the final distribution. Since the API only provides the log probabilities of the 20 most likely tokens at every token position, if one of Yes and No tokens do not appear, we consider its probability to be 0. If both do not appear, we consider their probabilities to be 0.5 each. We also note that we only consider the log probabilities of the token in the first position, as we expect a yes-no answer from GPT. To further encourage this, we make a small addition to the VQA prompt, concatenating “(yes/no)” to the end of the question.
We also ran into a couple issues with the Azure OpenAI API. Some specific examples in our dataset triggered Azure’s content filter, even when setting the filter’s threshold to high. To minimize this, we replaced “executed” with “complete” in the VQG prompt. In the cases where the issue persisted, we were forced to skip over the example during evaluation. Also, a small portion of API responses returned no completion content during VQG or question rephrasing for NLI (i.e. the content attribute of the response had a value of None). To deal with this, we implemented a retry procedure, where if no content is given by GPT, we prompt it again with the same prompt. If this second chance also provides no content, we either ignore the example during evaluation in the case of VQG, or simply concatenate the question and answer for question rephrasing. On the validation data, 28 out of 500 examples were omitted, and 103 question candidates could not be rephrased by GPT-4o.
The selected inference hyperparameters $n$, $\delta$, $\epsilon$, and $\tau$ are tuned as in our open-source model results, and listed in Table 10. The results on the validation data are listed in Table 11. Comparing GPT-4o to base VLMs we evaluated, it is generally outperformed open-source VLMs in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best models, though, outperform GPT-4o under all evaluation metrics.
$n$ $\delta$ $\epsilon$ $\tau$
4 0.4 0.1 0.19
Table 10: Selected inference hyperparameters for GPT-4o.
Partition Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 58.7 54.3 17.5 1.83 0.730
Table 11: Ego4D-PMD validation set results for GPT-4o.
E.7 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs shown in Section 4.4, in Figure 7, we provide several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.
Correct and coherent points.
Cyan points have low error with high informativeness and relevance, indicating correct decisions with coherent explanations. These are the best case examples from the model. Figure 7, Example A is one such case, where LLaVA correctly determines that the procedure “Pick up a sink brush from the kitchen slab” has been successfully completed, explaining it coherently and succinctly with a single question and answer about the location of the sink brush.
Incorrect and incoherent points.
Conversely, red to magenta points have high error, low informativeness, and low relevance, indicating incorrect decisions with incoherent explanations. These are the worst case examples from the model. Figure 7, Example B is one such case, where LLaVA incorrectly decides that the procedure “Grind the steel pole with the angle grinder” was not successfully completed due to the person in the image not wearing protective gear, an incoherent explanation for the decision.
Correct but incoherent points.
Indigo to black points have low error, but low relevance and informativeness, indicating correct decisions without sufficient explanation. Figure 7, Example C, is an instance of this, where LLaVA correctly decides that the person in the image has not successfully completed the procedure “Pick onion ring from the bowl of flour mixture with the fork,” explains it only by generating an answering a question about the location of a bowl of flour (an insufficient explanation).
Coherent but incorrect points.
White points have high error, relevance, and informativeness, indicating coherent explanations that do not lead to a correct decision. In other words, the information collected by the VLM should have been sufficient to make a correct decision (according to our automated coherence metrics), but this did not occur. Figure 7, Example D shows one such case, where LLaVA incorrectly decides that the procedure “Drop the bottle of mustard on the countertop” was unsuccessful. While it correctly identified that the bottle is on the countertop, which suggests the success of the procedure, it later mistakenly identified the bottle to be on the floor, causing it to make the wrong decision. The ability of this analysis to easily identify issues like this may be useful for future work in PMD and task guidance, as it enables the detection and thus the correction of bugs in the system’s reasoning.
Irrelevant but informative points.
Blue points have low relevance but relatively high informativeness, indicating irrelevant questions that still yield informative answers. As shown in Figure 7, Example E, this does not necessarily indicate a failure of LLaVA, rather a terse explanation. In this example, LLaVA correctly determines that the procedure “Pick the manual paper” has not been completed successfully. It reasonably explains this decision by asking about the presence of a manual paper and responding with No. The question of whether the person is working with a manual paper is deemed irrelevant by our metrics because if the answer were instead Yes, this would not provide sufficient information to conclude that the procedure was successful. However, since the answer was No, we do have sufficient information to conclude that the procedure is unsuccessful, despite the question being relatively indirect. Blue points may thus point to sufficient explanations which lack some detail or specificity.
Relevant but uninformative points.
Green and yellow points have high relevance but low informativeness, indicating a failure to extract useful information in VQA. Green points have close to zero informativeness, indicating unsure responses in VQA. In Figure 7, Example F, LLaVA explains its decision about the procedure “Stir the chocolate inside the blender” by asking whether the chocolate and the spoon are in the blender in various ways. However, these objects are not present in the image and thus LLaVA’s answer is not confident, causing it to respond Unsure to most questions, leading to low informativeness. Despite its failure to answer questions, LLaVA still arrives at the correct conclusion that the procedure has not been successfully completed.
Meanwhile, yellow points have highly negative informativeness, indicating counterproductive responses in VQA that oppose the correct decision. As shown in Figure 7, Examples G and H, these cases typically occur when the VLM does not recognize an object in the image, or it recognizes an object that is not in the image. In Example G, LLaVA incorrectly decides that the procedure “Put the trowel in a bin” is unsuccessful because it does not recognize that the trowel is indeed in a bin, perhaps because it is relatively small in the image and does not contrast from the background. In Example H, LLaVA incorrectly decides that the procedure “Put the bottle in the cabinet” is successful because it hallucinates that a bottle is in the cabinet, despite neither object appearing in the image. The ability of this analysis to easily identify failures of visual perception in VLMs again may be useful for future work in this area.
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label: ✓        Predicted: ✓ Explanation: 1. Is the sink brush in the person’s hands? Yes
Example B: Grind the steel pole with the angle grinder. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the person wearing safety goggles? No 2. Is the person wearing any protective gear? No 3. Is the person wearing gloves? No
Example C: Pick onion ring from the bowl of flour mixture with the fork. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the bowl of flour mixture on the table? Yes
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the bottle of mustard on the countertop? Yes 2. Is the bottle of mustard on the floor? Yes
Example E: Pick the manual paper. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the person working on a manual paper? No
Example F: Stir the chocolate inside the blender with the spoon. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the chocolate in the blender? Unsure 2. Is the chocolate inside the blender? Yes 3. Is the spoon in the blender? Unsure 4. Is the spoon in the chocolate? Unsure
Example G: Put the trowel in a bin. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the trowel in a bin? No
Example H: Put the bottle in the cabinet. Refer to caption Label: ✗        Predicted: ✓ Explanation: 1. Is the bottle in the cabinet? Yes
Figure 7: Sample outputs from LLaVA with coherence-based ranking, representing the range of behaviors visualized in Figure 4. Some images cropped for clarity and space.