### v2

# Transparent and Coherent Procedural Mistake Detection

Abstract
Procedural mistake detection (PMD) is a challenging problem of classifying whether a human user (observed through egocentric video) has successfully executed a task (specified by a procedural text). Despite significant recent efforts, machine performance in the wild remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we extend PMD to require generating visual self-dialog rationales to inform decisions. Given the impressive, mature image understanding capabilities observed in recent vision-and-language models (VLMs), we curate a suitable benchmark dataset for PMD based on individual frames. As our reformulation enables unprecedented transparency, we leverage a natural language inference (NLI) model to formulate two automated metrics for the coherence of generated rationales. We establish baselines for this reframed task, showing that while VLMs struggle off-the-shelf, their accuracy, coherence, and efficiency can be improved by incorporating these metrics into common inference and fine-tuning methods- though not without tradeoff. Lastly, our multi-faceted metrics visualize common outcomes, highlighting areas for further improvement.
Transparent and Coherent Procedural Mistake Detection
Shane Storks        Itamar Bar-Yossef        Yayuan Li        Zheyuan Zhang        Jason J. Corso        Joyce Chai University of Michigan, Ann Arbor, Michigan, USA {sstorks, itamarby, yayuanli, zheyuan, jjcorso, chaijy}@umich.edu
Figure 1: To reason through the complex task of procedural mistake detection (PMD), vision-and-language models (VLMs) are conditioned to gather visual evidence through an iterative self-dialog to rationalize their final decision.
1 Introduction
The problem of interactive task guidance has recently attracted attention in the AI research community Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and interact with them through language to guide them to complete a task. A key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must anticipate the success conditions for actions in this procedural text, then verify them by extracting relevant physical state information from the visual scene.
Prior work in PMD has explored a mix of specialized (primarily vision-based) classifiers Sener et al. (2022); Wang et al. (2023a); Peddi et al. (2024) as well as foundational language models (LMs) and vision-and-language models (VLMs) Du et al. (2023); Bao et al. (2023); Peddi et al. (2024); Flaborea et al. (2024), but this problem has proven difficult, and current approaches fail to achieve accuracy above chance in zero-shot PMD in the wild. Qualitatively, Bao et al. (2023) find that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are unreliable mistake detectors, transparency is crucial for the user to understand the system’s reasoning process and act on or disregard it accordingly. and continued improvement.
To promote the development of PMD systems with transparent and justified decisions, we propose a reformulated problem of coherent PMD: given a procedural text and egocentric video frame, VLMs must not only classify whether a mistake has occurred, but also support this decision with a rationale consisting of evidence from the visual scene. As shown in Figure 1, this rationale takes the form of an iterative self-dialog of generated questions and yes-no answers which directly condition classification. [^2] This choice of rationale resembles visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), long-studied multimodal tasks that VLMs excel at Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). Since recent VLMs struggle to extract detailed information from videos, but have exhibited more mature image understanding capabilities, we curate an approachable large-scale dataset for PMD based on individual video frames annotated in Ego4D Grauman et al. (2022). We define two metrics for the coherence of generated rationales based on a natural language inference (NLI) model. To lay a foundation for research in coherent PMD, we establish baselines by exploring three natural interventions to VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness VLMs’ in-context learning capability to generate additional candidate questions based on human-written examples, and (3) we use our metrics to fine-tune VLMs to generate more coherent questions. Our results show that while VLMs struggle off-the-shelf, these interventions can improve VLMs’ accuracy, coherence, and rationale generation efficiency, albeit creating tradeoffs between these properties. We lastly show how our multi-faceted metrics visualize common outcomes in coherent PMD, enabling fine-grained evaluation and identification of areas for future improvement.
Figure 2: Selected examples from Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD). For each matching pair of a video frame and procedural text, we generate a success example and various mistake examples by sampling alternate video frames: incomplete execution, execution with the wrong verb (e.g., wringing a cloth instead of folding), execution with the wrong noun (e.g., folding paper instead of a cloth), and execution with both the wrong verb and noun (e.g., opening a notepad instead of folding a cloth). Images cropped for space.
2 Problem Formulation and Dataset
In this section, we define the extended problem of coherent PMD in an approachable manner for VLMs, describe how to apply VLMs to the problem, then lastly introduce a benchmark dataset we curated for evaluating coherent PMD.
2.1 Defining Coherent PMD
The inputs for PMD are a short procedural text $T$ and a single video frame $F$, which may or may not visualize the successful completion of the procedure described in $T$. Given these inputs, a system should return a binary success decision $y$ for whether the procedure has been successfully completed ($y=0$ indicates success, and $y=1$ indicates the detection of a mistake). In coherent PMD, it must additionally generate a rationale $R=(\mathcal{Q},\mathcal{A})$, where $\mathcal{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$ and $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$ are respectively sequences of $n$ yes-no questions and their predicted answers, which provide evidence for the decision.
2.2 Applying VLMs to Coherent PMD
As shown in Figure 1, we elicit a rationale from VLMs through a self-generated visual dialog Das et al. (2017) consisting of visual question generation (VQG) based on the procedural text and dialog history, and visual question answering (VQA) based on the video frame. The evidence compiled in this rationale then conditions a success classification for whether the procedure has been completed in the given video frame. This structure goes beyond past approaches for PMD using VLMs; while Du et al. (2023) only elicited success classification, Bao et al. (2023) used procedure-agnostic prompts to caption images before classification, nonetheless disregarding this information in quantitative evaluations.
VQG.
We prompt the VLM to generate a series of questions given the procedural text (and previous questions and answers in later iterations, enabling deductive reasoning). To encourage logical, diverse questions, we apply greedy beam search to generate 4 candidates, from which we select the most likely candidate not generated previously.
VQA.
After a question is generated, it is answered by the VLM given only the question and video frame. If the probability of the chosen answer (i.e., “Yes” or “No”) exceeds an answer sureness threshold of 60%, we append it to the dialog history, otherwise we append “Unsure.” [^3] Unsure answers are excluded from example-level informativeness (defined in Section 3), and excluded from previous questions and answers in metric calculations.
Success classification.
After each iteration of VQG and VQA, we prompt the VLM to judge whether the procedure has been successfully executed based on the video frame and entire dialog history. The VLM’s decision is made using a mistake confidence threshold $\tau$ (tuned on the validation data for each approach) on its mistake likelihood. The prompt and answer for this step are excluded from the dialog history in future iterations.
Stopping criteria.
To prevent over-generating evidence, which could introduce noise and degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before $n$ questions are generated) if the success likelihood stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). $n$ is fixed at 10, while $\delta$ and $\epsilon$ are tuned based on the validation data for each presented approach.
2.3 Constructing a Dataset for PMD
We follow Du et al. (2023) in recasting Ego4D Grauman et al. (2022), a large-scale egocentric video dataset for everyday activities with dense annotations for various aspects of the videos, into an offline mistake detection format, but expand the diversity of mistake types studied there. Ego4D’s hand and object interactions data subset includes videos of physical actions being performed with various objects. Each video is annotated with narrations describing fine-grained procedures being performed, timestamps for when it begins and ends, and category labels for the verb and noun characterizing the procedure. This makes an ideal testbed for evaluating VLMs’ understanding of real-world actions in video frames, but the data is not formulated for mistake detection. We thus apply several preprocessing steps to the data to create a new Ego4D for Procedural Mistake Detection (Ego4D-PMD) benchmark that includes successful cases and a breadth of mistake types for each annotated procedure, visualized in Figure 2.
Additional details about these preprocessing steps and summary statistics for the Ego4D-PMD dataset are presented in Appendix A. As listed there, to conserve compute, we randomly sample a subset of 10,000, 500, and 2,000 examples respectively from the training, validation, and testing partitions (evenly split between success and mistake cases) for the forthcoming experiments.
3 Evaluating Coherence in PMD
Next, we describe our application of a fine-tuned NLI model to calculate two evaluation metrics for coherent PMD: relevance of questions and informativeness of answers to those questions.
3.1 Using NLI Models to Judge Success
As shown in Figure 3, LMs fine-tuned for NLI can estimate the sufficiency of visual questions and answers in rationalizing whether a procedure was successfully completed. This requires an NLI model $f_{e}$ (which returns a probability that an input premise string entails a hypothesis string), a premise transformation $t_{p}$ (which converts a question $Q$ and answer $A$ into a declarative statement to add to the premise), and a hypothesis prompt template $t_{h}$ (which creates a hypothesis about the success of the procedure in $T$). We then calculate the NLI model’s probability for the success of procedure $T$ based on the rationale $R=(\mathcal{Q},\mathcal{A})$:
$p_{e}(T|\mathcal{Q},\mathcal{A})=f_{e}\bigl{(}t_{h}(T)|\{t_{p}(Q_{i},A_{i})|1\leq i\leq n\}\bigr{)}$
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^4] See https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^5] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^6] A template was used as we found complex procedural texts were unlikely to be rephrased accurately, degrading correlation with human judgments (see Appendix C).
Figure 3: Using BART Lewis et al. (2020) fine-tuned on MNLI Williams et al. (2017) to judge success.
3.2 Relevance
A coherent decision in PMD should be supported by relevant questions about the state of the environment. [^7] For example, given a procedure “In a bowl, add the cut cherry tomatoes” Peddi et al. (2024), the question “Are there tomatoes in the bowl?” is relevant to the success of the procedure, while the question “Is the bowl blue?” is not. We measure the relevance of a question $Q^{\prime}$ to the success of a procedure $T$, given previous questions $\mathcal{Q}$ and their answers $\mathcal{A}$, as follows:
$\begin{split}\texttt{Rel}(Q^{\prime}|T,\mathcal{Q},\mathcal{A})=&\bigl{|}p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup\text{``{No}''})\\ &-p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup\text{``{Yes}''})\bigr{|}\end{split}$
This definition quantifies how much impact the answer to the proposed question $Q^{\prime}$ can have on the success probability (as estimated by the NLI model). If the success probability is similar for “Yes” and “No” answers, this suggests that $Q^{\prime}$ would not reveal pertinent information (i.e., beyond that in $\mathcal{Q}$ and $\mathcal{A}$) about whether the procedure in $T$ was successfully executed by the user, and thus the relevance would be low. If the success probabilities vary widely depending on the answer, this suggests that $Q^{\prime}$ can reveal important new information to help make the decision.
Example-level relevance.
To reward systems that propose consistently relevant questions in our evaluations, we summarize the relevance of a sequence of questions generated for a particular example by taking the mean question relevance with respect to previous questions and answers:
$\frac{1}{n}\sum\limits_{i=1}^{n}\texttt{Rel}\bigl{(}Q_{i}|T,\{Q_{j}:j<i\},\{A_{j}:j<i\}\bigr{)}$
3.3 Informativeness
Since a relevant question does not guarantee an informative answer, [^8] For example, in the procedure “In a bowl, add the cut cherry tomatoes,” “Are there tomatoes in the bowl?” is a relevant question, but a “Yes” answer is insufficient to confirm 100% completion (more tomatoes could be outside the bowl). and VLM errors in answering questions could unintentionally introduce conflicting information, it is necessary to evaluate the sufficiency of VLMs’ answers in justifying a PMD decision. To achieve this, we measure the informativeness of a predicted answer $A^{\prime}$ for a question $Q^{\prime}$ to the success of a procedure $T$ (given previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$) as follows:
$\texttt{Inf}(A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A})=1-H\bigl{(}p_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup A^{\prime})\bigr{)}$
$H$ is the binary Shannon entropy of the success probability $p_{e}$, calculated by $H(p)=-p\log_{2}p-(1-p)\log_{2}(1-p),p\in[0,1]$. This definition for informativeness quantifies how much information an answer to a question provides in determining the success of the procedure. As such, if the success probability given this answer $A^{\prime}$ to $Q^{\prime}$ is confident (i.e., very low or high), this indicates that $A^{\prime}$ (along with $\mathcal{Q}$ and $\mathcal{A}$) are sufficient to make a decision, and thus informativeness is high. Conversely, a success probability close to 50% suggests the evidence gathered thus far is insufficient, yielding low informativeness. Informativeness is expressed as a number of bits between 0 and 1.
Reference-adjusted informativeness.
We also wish to account for cases where the evidence gathered in questions and answers indicates the wrong PMD decision. To do so, we define the NLI model’s PMD belief $y_{e}(T|\mathcal{Q},\mathcal{A})$ as 1 (mistake) if $p_{e}(T|\mathcal{Q},\mathcal{A})<0.5$, else 0 (success). Given the ground truth PMD label $y^{*}$, we then define the reference-adjusted informativeness to be negative if the NLI model judges the evidence gathered as indicating the wrong success decision:
$\texttt{Inf}^{*}(A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A},y^{*})=\\ \begin{cases}\texttt{Inf}\bigl{(}A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A}\bigr{)},&y_{e}(T|\mathcal{Q}\cup Q^{\prime},\mathcal{A}\cup A^{\prime})=y^{*}\\ -\texttt{Inf}\bigl{(}A^{\prime}|Q^{\prime},T,\mathcal{Q},\mathcal{A}\bigr{)},&\text{else}\end{cases}$
Example-level informativeness.
To summarize the sufficiency of information gathered throughout the self-dialog (which may have some uninformative answers), we take the maximum (reference-adjusted) informativeness across the self-dialog:
$\max\limits_{1\leq i\leq n}\texttt{Inf${}^{*}$}(A_{i}|Q_{i},T,\{Q_{j}:j<i\},\{A_{j}:j<i\},y^{*})$
4 Rationale Coherence Interventions
To validate these metrics and examine the relationship between rationale coherence, PMD accuracy, and rationale generation efficiency, we next introduce two inference-time interventions to encourage VLMs to select more coherent questions from generated candidates, as well as a preference optimization approach to encourage generating more coherent candidates based on coherence metrics.
4.1 Coherent Question Selection
While a typical beam search would use a sequence likelihood-based approach to rank candidate questions, an alternative is to re-rank candidates using the reference-free coherence metrics introduced in Section 3. This could encourage selecting questions that are likely to bring in new, salient, and helpful information. Furthermore, as adaptive information-seeking is a core component of humans’ reasoning capabilities Coenen et al. (2019), we will supplement candidate questions with candidates generated through in-context learning from human-written examples (which we assume are reasonably coherent and effective).
Next, we introduce these two approaches we use to augment the candidate question pool for more coherent candidates: coherence-based re-ranking and candidate generation through in-context learning. We then compare their performance on Ego4D-PMD with that of vanilla VLMs.
4.1.1 Coherence-Based Question Selection
We implement a coherence-based candidate question re-ranking approach as follows. Given a set of question candidates $\hat{\mathcal{Q}}$ for procedural text $T$ along with previous confidence-filtered questions $\mathcal{Q}$ and answers $\mathcal{A}$, we can select the best question $Q^{*}$ by maximizing the product of relevance and potential informativeness across all $Q\in\hat{Q}$:
$\texttt{Rel}(Q|T,\mathcal{Q},\mathcal{A})\mathop{\times\>\>\text{max}\>\>\texttt{Inf}}_{A\in\{\text{{Yes}},\text{{No}}\}}(A|Q,T,\mathcal{Q},\mathcal{A})$
This ranking prioritizes well-rounded questions that can yield impactful information for success classification which leads to the most confidence. $Q^{*}$ is then concatenated to the dialog history and answered by the VLM as described earlier.
4.1.2 In-Context Learning Augmentation
Applying in-context learning in coherent PMD is not straightforward, as it would require reasoning over multiple images and dialogs of varying length. Instead, we use in-context learning to improve the text-based VQG step by providing examples of human-written questions. We achieve this by annotating 20 procedures from the Ego4D training data with 3 reasonable questions one could ask about a given procedure to judge its success. [^9] All questions and more details in Appendix E.1. We prompt the VLM with these example procedures and questions, the current procedure at hand, and the previous 2 questions proposed by the VLM (as available) to incorporate information the VLM already collected. We then generate 4 additional candidate questions using the same constraints described earlier. To minimize the impact of ordering, in-context examples are randomly shuffled in every prompt. [^10] Examples are equivalently shuffled in compared results.
4.2 Coherent Question Generation
While the above approaches may boost the coherence of PMD rationales, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. As such, we also explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
Specifically, we apply direct preference optimization (DPO; Rafailov et al., 2023) with quantized low-rank adaptation (QLoRA; Hu et al., 2022; Dettmers et al., 2024) to fine-tune a VQG adapter for LLaVA. We generate training data by first running inference over the Ego4D-PMD training data with coherence-based candidate question re-ranking and additional candidates generated through in-context learning. [^11] Ablation without in-context learning in Appendix E.2. We then fine-tune VLMs on pairs of chosen and rejected candidate questions for each self-dialog turn based on their coherence ranking score. [^12] More details and validation results in Appendix E.3. The top ranked question is always chosen, while a rejected question is sampled from the bottom half of candidates. At inference time, the trained adapter is applied during VQG, and disabled for other steps.
5 Experimental Results
In this section, we evaluate VLMs on coherent PMD off-the-shelf and under the previously introduced interventions. Metrics include mistake detection accuracy, mean example relevance and reference-adjusted informativeness of rationales (defined in Section 3). We add two metrics for self-dialog efficiency: the average number of iterations that occurred before stopping (i.e., the length of the dialog), and the average information gain in the success likelihood across all iterations in bits (i.e., how much information the VLM got from it). We lastly show that our proposed metrics visualize common behaviors of VLMs, enabling a panoptic understanding of their performance.
We specifically evaluate InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024), small open-source VLMs feasible for online use (important for real-world applications like task guidance). They apply different architectures and training strategies to integrate vision into their LMs. LLaVA and InstructBLIP were not trained on Ego4D, while Llama 3’s training data was not disclosed. For a reference point with state-of-the-art proprietary VLMs, we additionally include results with off-the-shelf GPT-4o OpenAI et al. (2024b). Additional details and prompt templates can be found in Appendix D and E.4.131313In Appendix E.5, we include an additional evaluation of VLMs without generating rationales (as done in prior work).
InstructBLIP
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 63.5 17.5 .224 2.84 .263
L 65.2 13.9 .340 4.71 .358
C 64.6 25.5 .281 3.46 .293
C 66.6 35.2 .359 3.47 .359
LLaVA
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 60.7 40.3 .259 3.25 .435
L 61.8 36.5 .272 3.34 .429
C 61.4 66.5 .321 3.06 .540
C 67.8 75.5 .464 3.46 .663
Llama 3
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 61.0 16.5 .275 4.70 .223
L 59.1 15.9 .317 6.51 .256
C 60.2 25.2 .341 6.35 .264
C 61.7 52.5 .436 3.59 .379
GPT-4o
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 55.4 54.0 .220 1.84 .793
Table 1: Ego4D-PMD test set results for GPT-4o OpenAI et al. (2024b) and likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
LLaVA + DPO
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.2 75.7 .318 2.33 .617
L 63.7 58.5 .330 2.67 .548
C 62.3 92.2 .340 2.06 .719
C 64.2 95.0 .304 1.81 .742
Table 2: Ego4D-PMD test set results for LLaVA with coherence-based fine-tuning through DPO.
5.1 Coherent Question Selection Results
Experimental results for coherent question selection interventions are presented in Table 2.141414Hyperparameters and validation results in Appendix E.6. We find that coherence-based re-ranking and in-context learning [^15] We visualize the distribution of question sources (i.e., in-context learning vs. full self-dialog history) in Appendix E.7. both sharply improve the relevance and informativeness in all models, reaching a respective 75.5% and 0.464 bits in LLaVA. This suggests that questions that VLMs find most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 67.8%, demonstrating that coherent rationales are valuable to accurate PMD. The information gain under our question selection interventions is consistently higher, reaching a maximum of 0.663 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent rationales. [^16] To better understand performance gains here, in Appendix E.8, we contextualize these results with a naïve semantic diversity-based ranking baseline. Meanwhile, our best configurations outperform GPT-4o in accuracy, relevance, and informativeness at less than 10% of its size Abacha et al. (2025). While GPT-4o takes fewer iterations and exhibits higher information gain (thus making faster and more confident decisions), its rigidity as a closed proprietary VLM prevents further improvements.
5.2 Coherent Question Generation Results
When fine-tuned for coherent question generation, Table 2 shows that relevance drastically increases from a previous maximum of 75.5% to 95.0%. Remarkably, this demonstrates that VLMs can learn to ask more coherent questions for PMD, a task based on properties of the physical world.
However, we find that accuracy and informativeness drop slightly from the best results in Table 2 to a respective maximum of 64.2% and 0.340 bits. This suggests that asking more relevant questions is not enough to improve performance globally. We suspect this is due to an implicit difficulty trade-off of asking highly relevant questions, which strongly indicate success if answered one way, otherwise a mistake. For elaborate procedures, highly relevant questions may cover multiple aspects or states of a scene. As such, answering these more complex questions may be difficult for VLMs. For example, while the VLM-generated question “Is the soil placed around the seedling with the trowel in the person’s hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering it requires understanding spatial relations between several objects (i.e., soil, seedling, trowel, and hand). Future work may explore methods to encourage VLMs to generate simpler questions. [^17] As an initial inquiry here, we experiment with applying an exponential length penalty to VQG in Appendix E.9.
Meanwhile, we find that fine-tuning reduces the average number of iterations taken to come to a conclusion by about 1.25, reaching a minimum of 1.81 (matching that of GPT-4o). This is especially important for online PMD, which requires a fast reaction. Further, the average information gain reaches a new peak of 0.742, nearly matching that of GPT-4o. This suggests that coherence-based fine-tuning can empower VLMs to make more confident decisions faster. We lastly observe that coherence-based ranking and in-context learning have a smaller impact on performance metrics in fine-tuned LLaVA than in the base model (Table 2). This suggests that fine-tuning VLMs using our coherence metrics reduces the need for inference-time interventions for coherence.
Vanilla LLaVA
+ Coherence-Based Ranking
+ In-Context Learning
+ Coherence-Based DPO
Figure 4: Visualization of decision error, relevance, and reference-adjusted informativeness for configurations of LLaVA on Ego4D-PMD testing examples. Each data point’s color indicates its position along each axis.
5.3 Visualizing PMD Performance
An additional benefit of our automated coherence metrics is the ability to audit the reasoning behaviors of VLMs. In Figure 4, we visualize the distribution of decision error, example-level relevance, and example-level informativeness of four representative approaches. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based fine-tuning with DPO. For each example, decision error is calculated by how far the VLM’s success likelihood is from being 100% confident in the ground truth success decision. Point colors indicate combinations of decision error, relevance, and informativeness, highlighting common local outcomes. Several examples are described in Appendix E.10.
In comparing the plots globally, we see a virtual elimination of red, black, and indigo points with irrelevant and uninformative rationales. We also see a significant reduction in the range of informativeness under coherence-based fine-tuning. This suggests that when questions are answered correctly under this approach, it is highly informative to the success decision, and vice versa. This is in line with our earlier observation that coherence-based fine-tuning may encourage the generation of overly relevant and thus complex questions, which can make or break the rationale. There are also a large number of points with zero informativeness, i.e., unsure answers to these complex questions.
6 Related Work
Beyond the prior work in PMD discussed in Section 1, we next discuss other related research.
6.1 Multi-Step Reasoning in LMs & VLMs
LMs have exhibited impressive reasoning capabilities from prompting methods Wei et al. (2022); Kojima et al. (2022), later strengthened with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024); Qi et al. (2024), and fine-tuning on reasoning chains from stronger LMs Wang et al. (2024); Gou et al. (2024); Muennighoff et al. (2025) and/or self-generated during reinforcement learning OpenAI et al. (2024a); DeepSeek-AI et al. (2025). Related to PMD, some work has investigated LMs’ reasoning about dependencies between physical procedures Bellos et al. (2024); Lal et al. (2024). Meanwhile, other work has applied iterative self-questioning approaches to deepen inquiry in other domains, e.g., medicine and fact verification Cattan et al. (2024); Li et al. (2025); Vladika et al. (2025).
In light of challenges in visual reasoning in VLMs Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024), some work has proposed training-free strategies and training paradigms to reduce visual hallucination Wan et al. (2024); Leng et al. (2024); An et al. (2024); Ganz et al. (2024), and utilized LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step You et al. (2023); Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024); Zhu et al. (2024); Sun et al. (2024); Jing and Rahman (2025). The latter stems from prior efforts in visual dialog Das et al. (2017); Kottur et al. (2019), which has been used to study and improve machine performance in various problem areas Chattopadhyay et al. (2017); de Vries et al. (2017); Uehara et al. (2022); Wang et al. (2022). Other works have assessed (V)LMs’ reasoning about procedures based on visual information Yang et al. (2018); Hendricks and Nematzadeh (2021); Jin et al. (2022); Yuksekgonul et al. (2023); Nguyen et al. (2024, 2025). In this work, we evaluated self-generated visual dialogs from VLMs in a novel problem of coherent PMD.
6.2 Leveraging NLI in Other NLP Tasks
NLI requires determining whether a premise text entails a hypothesis text, a reasoning challenge long studied in NLP Dagan et al. (2005). Many human-annotated resources have been compiled for NLI, thus significant progress has occurred in developing NLI models Storks et al. (2019). Prior work has leveraged this progress by using NLI models to improve the competence, confidence, and coherence of LMs in other tasks, such as dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), VQA Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). Building on this, we adopt two NLI-based metrics to strengthen VLMs in the novel problem of coherent PMD.
7 Conclusion
In this work, we evaluated foundational VLMs on a challenging problem of coherent PMD, where visual questions and answers must be generated to drive success decisions. To evaluate these rationales, we leveraged an NLI model to define two coherence metrics, using them to encourage coherent question selection and generation through common interventions. Our results showed that VLMs do not generate coherent rationales off-the-shelf, but these interventions improve their coherence, with the former also improving accuracy, and the latter improving the efficiency of generating and extracting information from rationales, albeit creating tradeoffs between these aspects. Further, patterns in accuracy and coherence metrics revealed detailed performance insights. Ultimately, when choosing an approach for coherent PMD, accuracy, coherence, and efficiency must be weighed for the setting. For example, while online applications may require shorter dialog (even at a cost of accuracy), high-risk applications may prioritize accuracy and confidence. This work lays a foundation for future research in the application of VLMs to PMD and task guidance, and investigation of how these system goals interact.
8 Limitations
Latency of self-dialog.
One limitation of our coherent PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings. This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.
Inherent limitations of single video frames.
Next, single video frames are limited in representing actions, which involve movement and state change. Our decision to focus on individual frames stemmed from preliminary experiments we performed with existing open-source VLMs for video understanding Lin et al. (2024); Li et al. (2023a), which are still in early stages. There, we found that they often confused information from frames in different regions of the video, preventing them from judging the final states of objects and reconciling this information with success of procedures, and resulting in poor performance. As such, our choice to focus on single images simplified the problem for current VLMs, enabling our experiments to begin building a meaningful understanding of their capabilities in PMD. To minimize the dependence on multiple frames in detecting mistakes, we applied several careful preprocessing steps to Ego4D-PMD (as discussed thoroughly in Section 2.3 and Appendix A). It is also worth noting that by not incorporating modalities beyond text and images, e.g., audio, the VLMs we studied are inherently limited in their capturing of physical information Yu et al. (2022); Zong et al. (2024). As the ability to reason over sequences of frames and other modalities evolves in state-of-the-art VLMs, future work can revisit this formulation and explore new approaches for them to reason over dynamic scenes.
Lack of ground truth rationales.
Lastly, it is worth noting that the dataset collected does not include ground truth rationales for mistake detection labels. Instead, we opted to propose automated coherence metrics for generated questions and answers based on a fine-tuned NLI model, which itself is prone to error and thus limits the objectivity of our evaluations. We chose this path for two important reasons. First, there may be multiple valid ways to detect a mistake through asking and answering visual questions, each of which could consist of different questions and/or different numbers of questions. [^18] For example, in trying to determine the success of the procedure “In a bowl, add the cut cherry tomatoes” from Peddi et al. (2024), we could reasonably ask one question “Are all the cherry tomatoes in the bowl?” or two questions “Are there cherry tomatoes in the bowl?” and “Are there any cherry tomatoes outside of the bowl?” In our opinion, existing metrics for text generation (which largely measure syntactic or semantic similarity of text) are not as well suited for this extremely challenging evaluation as fine-tuned NLI models (which are optimized to judge the logical consistency between texts). Second, in a real-world setting like PMD for task guidance, we believe that automated metrics are better suited for continually understanding and improving a deployed system than an offline benchmark of ground truth rationales. Despite this limitation, we believe that the value of these metrics is demonstrated by the accuracy and efficiency improvements brought by incorporating them into coherent PMD.
Acknowledgments
This work was supported in part by the DARPA PTG program HR00112220003. We would like to thank Megan Su, Ruixuan Deng, Fengyuan Hu, Andy Chung, Lu Wang, Richard L. Lewis, and the entire MSRP team for their helpful discussions, feedback, and insights.
Appendix A Ego4D-PMD Data Curation Details
Various benchmark datasets have been created for PMD from egocentric video, each of which includes video and procedural text along with various other modalities, as well as detailed information about mistakes Du et al. (2023); Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024). While these datasets are useful resources for research in task guidance, they are video-based and cover limited domains. Further, most of them include dialog interaction between a user and instructor agent which often causes mistakes to be corrected before or while they happen. They include mistakes around temperature, timing, small measurements, and other physical properties of the environment that are difficult for open-source VLMs, which are mostly optimized for representing single images, to perceive. While some of these issues could be overcome through a two-way dialog between the agent and a user, this makes it harder to isolate mistakes occurring in the videos and dive deep into the reasoning behind detecting them.
To alleviate these challenges and focus our inquiry, we follow Du et al. (2023) in recasting Ego4D, a procedural video dataset with a breadth of annotated and narrated everyday actions, into a single frame mistake detection format. The collection of the Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) data consisted of several preprocessing steps, outlined below.
Generating success examples.
As discussed above, Ego4D’s hand object interaction data is annotated in units of egocentric video clips of individual actions being performed by humans. We can form an example of a successful execution of the procedure by pairing each video clip with its annotated natural language narration of the procedure. Since most VLMs are not optimized to reason over multiple frames and videos, and those that are are still in very early stages, we sample exactly one frame from each video clip. Specifically, as each clip is carefully annotated with a postcondition time for the action, i.e., the time that the action has been completed, we simply sample the video frame at this annotated time and pair it with the text narration.
Generating mistake examples for incomplete procedures.
One natural type of mistake a user could make is not finishing a procedure. In addition to postcondition times, each video clip is annotated with a precondition time. Following a similar approach in Du et al. (2023), we can generate a mistake example by sampling a frame at the precondition time and pairing it with the video clip’s narration text. We expect that by doing this, the sampled frame will show the procedure at an incomplete state, and contain most of the same objects as the success example for the same clip. This poses a difficult challenge of identifying the key physical properties of the scene that would indicate completion.
Generating mistake examples for mismatched verbs and nouns.
Mistakes also happen when a user applies the wrong type of action to an object, causing an unexpected state, as well as when a user uses the wrong object or ingredient in a procedure. Following this intuition, we generate additional easier mistake examples from each clip by matching each clip with other clips that have a mismatched verb, noun, or both. While each clip is annotated with verb and noun categories, these categories are coarse-grained, making it impossible to guarantee that two clips with the same verb or noun label actually involve the same verb or noun, thus preventing sampling clips that share the same verb or noun. Instead, we apply the AllenNLP [^19] https://allenai.org/allennlp semantic role labeler to each narration text to identify the key participants in each procedure. For each clip, we then attempt to sample the postcondition frames from three mismatched clips: one with a mismatched verb (but matching nouns), one with a mismatched noun (but matching verb), and one with a mismatched verb and noun. We then pair these frames with the source clip’s narration text, creating mistake examples with varying levels of overlap with the source clip. While it is not always possible to find every such alternative clip for each clip in Ego4D, we can usually find at least one of them.
Transforming narrations into instructions.
The narration texts annotated in Ego4D are declarative statements about the actions being performed in each clip. This is not an accurate depiction of typical interactive task guidance and PMD settings, which usually revolve around instructional texts like recipes or guidebooks. As such, we convert each narration, e.g., “Someone washes the lettuce,” into imperative form, e.g., “Wash the lettuce,” using spaCy. [^20] https://spacy.io/ Further, some narrations describe procedures that are not suited for comparing physical state changes in text and images, such as social interactions, interactions with animals, interactions with electronic devices, and movements that are impossible to precisely characterize from the narration text (e.g., in “Move plate”). We use the verb and noun category annotations on each clip to filter out such cases.
Ensuring data quality.
We perform several additional steps to ensure high-quality mistake detection task instances. First, we remove clips where the precondition and postcondition frames are overly similar (i.e., at least 0.95 cosine similarity). We remove clips that are too dark (i.e., where the mean of all normalized RGB values is less than 0.2). When sampling frames from source clips, we sample several candidates within a small range around the precondition or postcondition timestamp, then select the least blurry candidate by the variance of the images’ Laplacian. Some videos in Ego4D show the same action being performed over and over (e.g., “Roll a ball of dough”), which can make it difficult to determine whether the state of the environment shown in a clip is the result of the current procedure or a prior one (given only a single frame). While future work applying video-optimized VLMs for coherent PMD in long-horizon tasks will need to address this challenge, this adds an unnecessary complexity to an already challenging task for current VLMs. As such, we remove any clips such that the same procedure in the clip has already been performed previously in the video. Lastly, we remove a handful of videos in Ego4D that we notice to be corrupted or significantly distorted.
Various statistics of the resulting full Ego4D-PMD dataset (and sub-samples used in the paper) are presented in Tables 3 and 4.
Type Train (Sample) Val. (Sample) Test (Sample) Total
Success 42.0k 5.00k 13.1k 250 18.1k 1.00k 73.1k
Mistake 99.4k 5.00k 25.4k 250 34,182 1.00k 159k
(Incomplete) 15.1k 755 4.91k 51 6.55k 194 26.5k
(Wrong V) 11.8k 604 2.69k 31 3.75k 108 18.2k
(Wrong N) 36.4k 1.85k 8.91k 87 11.8k 344 57.2k
(Wrong V&N) 36.1k 1.79k 8.91k 81 12,047 354 57.1k
Table 3: Distribution of example types in each partition of our curated Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD) dataset.
Type Train (Sample) Val. (Sample) Test (Sample) All
Verbs 83 80 77 55 78 71 83
Nouns 440 372 365 151 390 257 487
V-N Pairs 3,976 2,050 2,185 326 2,658 833 5,363
Table 4: Distribution of unique verb, noun, and verb-noun pair categories in each partition of our curated Ego4D-PMD dataset. Verb and noun categories are annotated for each narration in the Ego4D dataset.
Appendix B Rephrasing Questions and Answers for Coherence Evaluation
As discussed in Section 3.1, we use a fine-tuned NLI model to judge the success of procedures given questions and answers. In order to convert questions and answers into declarative statements to pass into the NLI model, we follow Srinivasan et al. (2024) in prompting a VLM with the following 10 in-context demonstrations of rephrasing before prompting it to rephrase a question and answer for the task at hand:
  1. 1.
    Question: Is there a bowl on the table?
    Answer: Yes
    Statement: There is a bowl on the table.
  2. 2.
    Question: Are the eggs cracked?
    Answer: No
    Statement: The eggs are not cracked.
  3. 3.
    Question: Does the cardboard box look open?
    Answer: Yes
    Statement: The cardboard box looks open.
  4. 4.
    Question: Are there any leaves outside of the basket?
    Answer: No
    Statement: There are not any leaves outside of the basket.
  5. 5.
    Question: Is the orange peeled?
    Answer: Yes
    Statement: The orange is peeled.
  6. 6.
    Question: Is the mug empty?
    Answer: No
    Statement: The mug is not empty.
  7. 7.
    Question: Are there hedge trimmers in the image?
    Answer: Yes
    Statement: There are hedge trimmers in the image.
  8. 8.
    Question: Has the light switch been turned on?
    Answer: No
    Statement: The light switch has not been turned on.
  9. 9.
    Question: Does the table have any cups on it?
    Answer: Yes
    Statement: The table has cups on it.
  10. 10.
    Question: Is the cabinet closed?
    Answer: No
    Statement: The cabinet is not closed.
Appendix C Coherence Metrics Human Study Details
In this appendix, we collect human judgments for the relevance and informativeness metrics defined in Section 3. To achieve this, we randomly sampled the outputs for 50 iterations of LLaVA’s self-dialog from two combinations of approaches:
  • Likelihood-based question ranking
  • Coherence-based question ranking augmented with question candidates from in-context learning
These outputs were from intermediate experiments, and thus the prompt used to initiate the self-dialog is slightly different than the one used for the experiments in the paper. However, as this is just a source of questions to compare human and machine judgments, this discrepancy does not impact the conclusions in the paper.
C.1 Annotation Instructions
For both relevance and informativeness annotation, we provided the following background for the task:
Imagine you just had eye surgery, and are currently unable to see. You’re performing a task you’re familiar with, but need help to determine whether you successfully completed it. You video call a friend (who is unfamiliar with the task) and show them what you’re working on. You then ask them some yes/no questions to figure out whether you successfully completed the task.
Relevance annotation instructions and example.
Annotators are provided the following instructions for annotating relevance:
For each annotation task, you will be given the following information:
  • A sentence describing the procedure you’re trying to perform.
  • An optional list of previous questions you already asked, and their answers.
  • A potential next question you could ask your friend.
You must rate how relevant the potential next question is. By relevant, we mean: given the previous questions and answers, how helpful could an answer to this question be in determining whether you successfully completed the task?
One sample is listed below:
Sentence: Drop the bowls on the table with your hand
Previous questions and answers:
  1. 1.
    Are the bowls on the table? (Answer: Yes)
  2. 2.
    Is the person holding the bowls in their hand? (Answer: No)
Potential next question: Is the person about to drop the bowls on the table?
Your rating:
  • 1 (very irrelevant)
  • 2 (slightly irrelevant)
  • 3 (neutral; may or may not be relevant)
  • 4 (slightly relevant)
  • 5 (very relevant)
  • Instructions Unclear
Informativeness annotation instructions and example.
Annotators are provided the following instructions for annotating informativeness:
For each annotation task, you will be given the following information:
  • A sentence describing the procedure you’re trying to perform.
  • A list of questions you asked your friend, and their answers.
You must rate how informative the questions and answers are. By informative, we mean: based on all the information you have, how sure are you about whether you succeeded?
One sample is listed below:
Sentence: Clean the bowl
Previous questions and answers: None
Last question: Is there a bowl in the image? Last answer: Yes
Your rating:
  • 1 (very uninformative/unsure)
  • 2 (slightly uninformative/unsure)
  • 3 (neutral; may or may not be relevant)
  • 4 (slightly informative/sure)
  • 5 (very informative/sure)
  • Instructions Unclear
C.2 Human Study Results
For each metric, we recruited 5 annotators (all English speakers with conferred or in-progress undergraduate degrees) to rate the relevance and informativeness of 10 machine-generated questions and answers. As this was an initial pilot study that was not scaled up further, annotators were authors of the work (not the first author) and/or peers of the authors, and not compensated. When comparing with our automated relevance and informativeness metrics, we use LLaVA for rephrasing.
Comparison of automated and human relevance judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., given the previous questions and answers, how helpful could an answer to this question be in determining whether the task was successfully completed) on a scale from 1-5 (least to most relevant). Between the resulting 50 annotations and corresponding automated metrics, we found a moderate Spearman correlation Spearman (1904) of $\rho=0.55$ ($p=0.000037$). This suggests that this automated measure of relevance is indeed positively correlated with human judgments of relevance.
Comparison of automated and human informativeness judgments.
We presented each annotator with 10 randomly selected VLM-generated questions $Q^{\prime}$ and answers $A^{\prime}$, along with previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$. Annotators were instructed to rate the relevance (i.e., based on all the information we have, how sure is the annotator about whether the procedure was successfully completed) on a scale from 1-5 (least to most informative). Between the resulting 50 annotations and corresponding automated metrics, we found a weaker Spearman correlation Spearman (1904) of $\rho=0.33$ ($p=0.020$). Interestingly, if we multiply the automated informativeness metrics by the relevance for $Q^{\prime}$, $\mathcal{Q}$, and $\mathcal{A}$, we find a stronger Spearman correlation of $\rho=0.50$ ($p=0.00022$). This suggests that while informativeness does have a relationship with human judgments, when multiplying it by relevance this relationship is stronger and more significant. This might be because the concepts of relevance and informativeness are themselves related. Intuitively, in most cases, a relevant question should be informative, and an irrelevant question should be uninformative. Meanwhile, fine-tuned NLI models could theoretically score an answer to an irrelevant question as informative, and vice versa. When proposing a coherence-based question re-ranking strategy in Section 4.1.1, we incorporated the inductive bias reflected in these human judgments by multiplying informativeness by relevance. Later, in Section 5.3, we visualized the full distribution of relevance and informativeness on our evaluation data to better understand this issue.
Appendix D VLM Self-Dialog Details
In this appendix, we include prompt templates and other supplementary details for the self-dialog we conditioned VLMs to generate, including visual question generation, visual question answering, and stopping criteria. To conserve GPU memory, VLM weights are 4-bit quantized at inference time, and are obtained from https://huggingface.co/Salesforce/instructblip-vicuna-7b, https://huggingface.co/liuhaotian/llava-v1.5-7b, and https://huggingface.co/meta-llama/Llama-3.2-11B-Vision.
D.1 Visual Question Generation
When prompting vision-and-language models (VLMs) to generate questions, we use the following prompt template:
[…]
Q:
Question generation is not conditioned on video frames, as we found significant performance degradation when VLMs were conditioned on the video frame while generating questions, often leading to completely nonsensical questions, e.g., “Is is is is is is?” To ensure the VLM generates yes-no questions, we constrain generation during decoding to enforce that each generated text begins with a word that can signal a yes-no question, [^21] Specifically, questions must begin with is, does, or has, along with all plural and past tense forms of these verbs. does not include the word or, and ends with a question mark. To avoid vague and high-level questions about the status of the procedure rather than low-level physical states, we prevent VLMs from repeating some words present in their input prompts: “successful,” “successfully,” “completed,” and “procedure.” To encourage logical questions while ensuring variety, we apply greedy beam search decoding with $k=8$ beams, returning the top 4 candidate questions to be ranked through LM likelihood or coherence metrics. [^22] Due to generation constraints, it is often the case that the VLM does not successfully generate all 8 candidates. In rare cases, these constraints prevent VLMs from generating any candidate questions; when this happens, we repeat the generation without any constraints (even though this may result in outputs that are not yes-no questions). Out of the candidates, we remove any that are exactly the same as previously generated questions, then select the most likely candidate based on the model’s log-likelihoods.
D.2 Visual Question Answering
Once a question is generated, we prompt the VLM with it and the video frame along with “A:” to elicit an answer. To produce the answer, we apply softmax over the output logits from the forward pass of the VLM for the Yes and No tokens. It is important to note that we exclude the dialog history from the context during VQA, as we again observed significant performance degradation when VLMs answered visual questions in the context of a longer dialog. This was especially prominent when several similar questions were generated and answered in a dialog, which often caused the VLM to creep from being initially unsure about the answer to being confidently wrong.
D.3 Success Classification
Questions and answers are generated iteratively until the stopping criteria described in Section 2.2 are met. To prompt the VLM to judge the success of a procedure after a question is answered, we use the following prompt:
[…]
Here, the logits of the Yes and No tokens are similarly used to produce a probability distribution over a success or mistake decision. The decision is determined by a mistake confidence threshold $\tau$, which is selected to maximize overall accuracy from a comprehensive set of 99 candidates $\tau\in\{0.01,0.02,\dots,0.98,0.99,1.0\}$.
D.4 Stopping Criteria Details
The stopping criteria hyperparameter $n$ is fixed at a value of 10. We chose to fix $n$ because tuning $n$ prevents objective comparison of the number of iterations taken by each approach, and $\delta$ and $\epsilon$ can still control the number of iterations of VLMs’ self-dialog. The hyperparameters $\delta$ and $\epsilon$ are selected from a grid search on the validation data over combinations of $\delta\in\{0.05,0.1,0.2,0.4\}$ and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a cascading summary metric which is the product of example-level informativeness (reference-adjusted) and example-level relevance if the VLM makes a correct mistake detection classification and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described in the previous subsection. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.6, while values for the results in Table 2 are listed in Appendix E.3.
Appendix E Supplementary Experimental Details and Results
In this appendix, we provide assorted details (e.g., prompt templates and hyperparameters) and supplementary results that were omitted from the main body of the paper.
E.1 Example Questions for In-Context Learning
As discussed in Section 4.1.2, we condition VLMs with sets of human-written questions for 20 procedures from the Ego4D for Procedural Mistake Detection (Ego4D-PMD) dataset. These human-written questions individually achieve 53.2% relevance and 82.1% maximum informativeness (i.e., for a yes or no answer) on average, with rephrasing done by LLaVA. The annotated procedures (underlined) and questions (italicized) are listed below:
  1. 1.
    Soak the sponge in a soapy water with your hands
    1. (a)
      Is there a sponge?
    2. (b)
      Is the sponge in water?
    3. (c)
      Is the water soapy?
  2. 2.
    Open the bottle
    1. (a)
      Is there a bottle in the image?
    2. (b)
      Is the bottle open?
    3. (c)
      Does the bottle have a lid on it?
  3. 3.
    Take the baking tray away from the table
    1. (a)
      Can you see a baking tray?
    2. (b)
      Is the baking tray on the table?
    3. (c)
      Is the baking tray picked up by someone?
  4. 4.
    Turn on a torch light
    1. (a)
      Is there a torch light in the photo?
    2. (b)
      Is the torch light powered on?
    3. (c)
      Is the torch light lit up?
  5. 5.
    Fold the right edge of the wrapper
    1. (a)
      Is there a wrapper in the image?
    2. (b)
      Is the wrapper completely flat?
    3. (c)
      Is the right edge of the wrapper folded?
  6. 6.
    Pour the water into the blue container
    1. (a)
      Do you see a blue container anywhere?
    2. (b)
      Is there water in the blue container?
    3. (c)
      Is the blue container empty?
  7. 7.
    Paint the patio with the paint brush
    1. (a)
      Is this a photo of a patio?
    2. (b)
      Is the patio painted?
    3. (c)
      Is someone holding a paint brush?
  8. 8.
    Spread the black peas on the salad with the spoon in your hand
    1. (a)
      Is there a salad?
    2. (b)
      Are there black peas on the salad?
    3. (c)
      Is there a spoon in someone’s hand?
  9. 9.
    Scoop paint from the pallet on the table with the paint brush
    1. (a)
      Do you see a paint brush and a paint palette?
    2. (b)
      Is there paint on the paint brush?
    3. (c)
      Is the paint brush in someone’s hand?
  10. 10.
    Wash the car with a sponge in your hand
    1. (a)
      Do you see a car?
    2. (b)
      Is the car clean?
    3. (c)
      Is the sponge being held?
  11. 11.
    Pick the scrubber from the sink
    1. (a)
      Do you see a scrubber somewhere?
    2. (b)
      Is the scrubber in the sink?
    3. (c)
      Is the scrubber in someone’s hand?
  12. 12.
    Peel the onion
    1. (a)
      Is there an onion in the image?
    2. (b)
      Is the onion’s skin removed?
    3. (c)
      Is the onion peeled?
  13. 13.
    Put the dirt in the dust bin
    1. (a)
      Is there a dust bin?
    2. (b)
      Is there dirt in the dust bin?
    3. (c)
      Is there any dirt outside of the dust bin?
  14. 14.
    Cut dough into two
    1. (a)
      Do you see any dough?
    2. (b)
      Is the dough in two pieces?
    3. (c)
      Is the dough whole?
  15. 15.
    Break the walnut with the nutcracker in your hand
    1. (a)
      Do you see a walnut?
    2. (b)
      Is the walnut cracked?
    3. (c)
      Is there a nut cracker in someone’s hand?
  16. 16.
    Turn off the tap
    1. (a)
      Is there a tap in the photo?
    2. (b)
      Is the water running?
    3. (c)
      Is the faucet switched off?
  17. 17.
    Heat the edge of the bag with the lighter
    1. (a)
      Do you see a bag and a lighter?
    2. (b)
      Is there a flame coming from the lighter?
    3. (c)
      Is the lighter near the bag?
  18. 18.
    Close the fridge
    1. (a)
      Is there a fridge?
    2. (b)
      Is the fridge open?
    3. (c)
      Can you see inside the fridge?
  19. 19.
    Chop green beans with a knife on the chopping board
    1. (a)
      Do you see green beans on a cutting board?
    2. (b)
      Are the green beans sliced?
    3. (c)
      Is someone using a knife?
  20. 20.
    Drop the brush in your hand on the oven
    1. (a)
      Is there a brush in the scene?
    2. (b)
      Is there an oven?
    3. (c)
      Is the brush on the oven?
E.2 DPO Training Data Composition Ablation
In fine-tuning VLMs to generate questions, we included questions generated through the in-context learning approach introduced in Section 4.1.2 to build upon our previous best result (i.e., augmenting VLMs with coherence-based ranking and in-context learning). In Table 5, we include an additional result for DPO where generated training data does not include candidate questions generated with in-context learning. As shown, compared to the results in Table 2, the best values of each metric are largely comparable with or without in-context learning, suggesting that in-context learning did not contribute significantly to the performance.
LLaVA + DPO (without ICL)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 61.0 84.3 .312 2.16 .746
L 64.8 64.9 .326 2.76 .622
C 61.2 94.6 .316 1.78 .772
C 61.8 93.0 .342 2.37 .790
Table 5: Ego4D-PMD test set results for DPO-trained VLMs without applying in-context learning in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.3 Question Generation Fine-Tuning and Inference Details
[…]
LLaVA + DPO
$n$ $\delta$ $\epsilon$
10 0.05 0.05
LLaVA + DPO (without ICL)
$n$ $\delta$ $\epsilon$
10 0.05 0.05
LLaVA + DPO (with length penalty)
$n$ $\delta$ $\epsilon$
10 0.05 0.05
Table 6: Selected training data generation hyperparameters for the results presented in Tables 2, 5, and 17.
LLaVA + DPO
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.5 10 0.2 0.05 0.78
L 10 0.2 0.05 0.75
C 10 0.2 0.05 0.58
C 10 0.1 0.1 0.76
LLaVA + DPO (without ICL)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 7.5e-6 0.5 10 0.1 0.05 0.59
L 10 0.1 0.05 0.85
C 10 0.4 0.05 0.66
C 10 0.05 0.05 0.64
LLaVA + DPO (with length penalty)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 7.5e-6 0.1 10 0.4 0.05 0.35
L 10 0.05 0.05 0.63
C 10 0.4 0.05 0.86
C 10 0.4 0.05 0.84
Table 7: Selected training and inference hyperparameters for the results presented in Tables 2, 5, and 17.
LLaVA + DPO
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 63.0 76.7 .328 2.27 .617
L 64.0 65.5 .347 2.48 .579
C 63.0 91.2 .342 2.11 .703
C 65.2 94.8 .318 1.85 .727
LLaVA + DPO (without ICL)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.0 85.0 .293 2.04 .757
L 62.6 71.9 .291 2.40 .689
C 62.6 93.6 .338 1.85 .756
C 63.0 92.0 .357 2.54 .783
LLaVA + DPO (with length penalty)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.0 93.1 .293 1.97 .739
L 65.4 66.1 .412 3.61 .648
C 61.6 97.1 .281 1.86 .756
C 61.8 97.2 .290 1.85 .758
Table 8: Ego4D-PMD validation set results for DPO-trained VLMs both with and without applying in-context learning in generating training data, as well as with a length penalty applied in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.4 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024b) [^23] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for coherent PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
To enable the GPT-4o evaluation, we make a few small changes to the VLM self-dialog implementation used in the main experiments. Since GPT can not be forced to generate yes-no questions with specific generation constraints, as done in our open-source implementation, we slightly modify the VQG prompt to encourage appropriate question generation. We add the following to the end of our original VQG prompt, with bold indicating the new text:
This is…from the photo. Generate an appropriate yes/no question.
Q:
To produce a probability distribution over the Yes and No tokens, used in both the VQA and success classification parts of the VLM self-dialog, we use the log probabilities provided by the API as logits are not directly available. We then normalize the probabilities to get the final distribution. Since the API only provides the log probabilities of the 20 most likely tokens at every token position, if one of Yes and No tokens do not appear, we consider its probability to be 0. If both do not appear, we consider their probabilities to be 0.5 each. We also note that we only consider the log probabilities of the token in the first position, as we expect a yes-no answer from GPT. To further encourage this, we make a small addition to the VQA prompt, concatenating “(yes/no)” to the end of the question.
We also ran into a couple issues with the Azure OpenAI API. Some specific examples in our dataset triggered Azure’s content filter, even when setting the filter’s threshold to high. In the cases where the issue persisted, we were forced to skip over the example during evaluation. Also, a small portion of API responses returned no completion content during VQG or question rephrasing for NLI (i.e. the content attribute of the response had a value of None). To deal with this, we implemented a retry procedure, where if no content is given by GPT, we prompt it again with the same prompt. If this second chance also provides no content, we either ignore the example during evaluation in the case of VQG, or simply concatenate the question and answer for question rephrasing. On the validation data, 28 out of 500 examples were omitted, and 103 question candidates could not be rephrased by GPT-4o.
The inference hyperparameters $n$, $\delta$, $\epsilon$, and $\tau$ are selected as in our open-source model results, and listed in Table 9. The results on the validation data are listed in Table 10, while the results on the testing data are listed in the main body of the paper in Table 2. Comparing GPT-4o to the base VLMs we evaluated, it is generally inferior in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best model configurations, though, outperform GPT-4o under all evaluation metrics except information gain. [^24] LLaVA with coherence-based fine-tuning, coherence-based re-ranking, and in-context learning achieves .742 bits of information gain, while GPT-4o achieves up to .793 bits. This shows that GPT-4o generally exhibits higher confidence despite having much lower PMD accuracy than our approaches, which is not necessarily an advantage. This suggests that GPT-4o is a reasonable starting point for coherent PMD, but like other off-the-shelf VLMs we evaluated, it may require additional interventions (e.g., coherence-based ranking) to be viable for this task.
$n$ $\delta$ $\epsilon$ $\tau$
10 0.4 0.1 0.19
Table 9: Selected inference hyperparameters for GPT-4o.
Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
58.7 54.3 .175 1.83 .730
Table 10: Ego4D-PMD validation set results for GPT-4o.
E.5 Rationale-Free Evaluation
For a reference point to incomparable prior works that have applied VLMs to PMD with a focus on classification accuracy Du et al. (2023); Bao et al. (2023), we additionally evaluate all studied VLMs from this work on a non-coherent PMD task. Specifically, we prompt each VLM with the following text:
[…]
This prompt is as comparable as possible to the one used for coherent PMD, but does not elicit a series of questions and answers from the VLM. We perform this evaluation on InstructBLIP, LLaVA, Llama 3, and GPT-4o. Table 11 lists the inference hyperparameters for this approach, while Table 12 lists the results for the validation and testing data. It is crucial to note that these results are not directly comparable to the coherent PMD results presented in the main body of the paper, as VLMs are not required to justify their decisions, removing the explainability enabled by coherent PMD (crucial for end users to interpret often incorrect VLM decisions and act on them accordingly). As prior work has already explored this setting more extensively, we do not intend to provide a rigorous study here, rather a reference point to compare how requiring the generation of rationales impacts PMD accuracy.
Interestingly, however, the rationale-free approach achieves generally better accuracy than the coherent PMD results in Tables 2 and 14 with likelihood-based ranking and no in-context learning. However, upon introducing coherence-based ranking and in-context learning, InstructBLIP and LLaVA achieve better accuracy than in the rationale-free approach. Further, the information gain in the rationale-free setting is consistently lower than those achieved in coherent PMD. This demonstrates that while the added transparency of incorporating rationales into PMD does cost the VLM some accuracy, improving the coherence of these rationales (e.g., through the approaches presented in this work) can recover this accuracy and more while enabling more confident decisions from VLMs. Rationale-free GPT-4o achieves the highest observed accuracy of 69.2%. Nonetheless, this accuracy is still low enough for errors to be common, thus necessitating the generation of a rationale for the user.
InstructBLIP LLaVA Llama 3 GPT-4o
0.34 0.77 0.54 0.98
Table 11: Selected values of mistake confidence threshold $\tau$ for rationale-free PMD with various VLMs. $n$, $\delta$, and $\epsilon$ are no longer used in rationale-free PMD, but $\tau$ is still tuned as in previous experiments.
InstructBLIP
Partition Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 62.6 0.00 .113
Testing 62.2 0.00 .117
LLaVA
Partition Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 64.4 0.00 .236
Testing 66.1 0.00 .233
Llama 3
Partition Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 65.8 0.00 .187
Testing 64.6 0.00 .176
GPT-4o
Partition Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 66.8 0.00 .718
Testing 69.2 0.00 .736
Table 12: Ego4D-PMD validation and testing set results for rationale-free PMD with various VLMs. As self-dialog rationales are no longer generated, relevance and informativeness cannot be calculated. Further, zero iterations are performed, and information gain is calculated using the average entropy of the success probability without any rationale.
E.6 Question Selection Inference Hyperparameters and Validation Results
In Table 13, we list the inference hyperparameters for the question selection results presented in Table 2: number of iterations $n$, early stopping $\delta$ and $\epsilon$, and mistake confidence threshold $\tau$. In Figure 5, we use detection error tradeoff (DET) curves to visualize the range of accuracy achieved with all candidate mistake confidence thresholds $\tau$ for the approaches compared in Table 2. The full validation set results with selected hyperparameters are listed in Table 14.
InstructBLIP
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 10 0.1 0.2 0.41
L 10 0.05 0.1 0.49
C 10 0.05 0.2 0.35
C 10 0.05 0.2 0.33
LLaVA
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 10 0.1 0.05 0.64
L 10 0.1 0.05 0.72
C 10 0.1 0.5 0.76
C 10 0.05 0.05 0.74
Llama 3
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
L 10 0.1 0.05 0.40
L 10 0.05 0.025 0.23
C 10 0.05 0.05 0.38
C 10 0.2 0.05 0.30
Table 13: Selected inference hyperparameters for the results presented in Table 2.
InstructBLIP
LLaVA
Llama 3
Figure 5: Mistake detection error tradeoff (DET) curves for VLMs applied to the Ego4D-PMD validation set with likelihood- and coherence-based candidate question selection approaches, with optional supplementary candidates generated through in-context learning (ICL).
InstructBLIP
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.0 18.3 .237 2.79 .265
L 61.8 14.0 .325 4.62 .358
C 63.8 26.0 .285 3.52 .298
C 64.2 36.2 .336 3.30 .363
LLaVA
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 62.0 42.7 .287 3.15 .437
L 62.2 41.7 .289 3.17 .439
C 63.6 68.5 .319 3.05 .532
C 64.4 76.5 .418 3.45 .659
Llama 3
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 60.8 16.8 .287 4.62 .219
L 61.2 16.2 .335 6.38 .245
C 64.2 25.0 .347 6.37 .236
C 64.8 52.3 .469 3.57 .396
Table 14: Ego4D-PMD validation set results for likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.7 Analysis of Question Sources in In-Context Learning
To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources on the validation data in Figure 6. As expected, candidates generated with in-context learning are relatively rarely selected under likelihood-based ranking, amounting to about 25.7% of VQG iterations for InstructBLIP, 9.0% of VQG iterations for LLaVA, and 5.2% of VQG iterations for Llama 3. On the other hand, they are selected more frequently in the coherence-based ranking, amounting to about 35.6% of VQG iterations for InstructBLIP, 30.6% of VQG iterations for LLaVA, and 36.5% of VQG iterations for Llama 3. Interestingly, in-context learning candidates are more dominant in earlier iterations, while candidates generated based on the dialog context are relatively more common in later iterations. This may suggest that after selecting a few questions from in-context learning in earlier iterations, the VLM is able to utilize them to generate better questions from the dialog context in later iterations. Alternatively, this could suggest that candidates from in-context learning have limited variety, and thus are less likely to be selected in later turns to avoid redundant questions or information.
InstructBLIP
LLaVA
Llama 3
Figure 6: Histograms of VLMs’ selected question sources, either self-dialog context or in-context learning (ICL) examples, by visual question generation (VQG) iteration for likelihood-based question selection (top) and coherence-based question selection (bottom).
E.8 Diversity-Based Ranking Baseline
One possible explanation for the effectiveness of coherence-based question ranking in Section 4.1 is that it enables the selection of more semantically diverse questions, thus collecting broader information about the image. To explore this question, we implement a supplemental diversity-based ranking approach which uses a sentence transformer Reimers and Gurevych (2020) [^25] See https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to embed all previous and candidate questions at each iteration, then selects the candidate question with the largest average cosine distance from previous questions.
As shown in Table 16, we observe that diversity-based ranking combined with in-context learning can also improve the accuracy of VLMs (to a level slightly below that of coherence-based ranking). Accuracy reaches respective maxima of 64.7% and 67.1% for InstructBLIP and LLaVA (compared to 66.6% and 67.8% under coherence-based ranking). This may suggest that some accuracy improvements in coherence-based ranking could be attributed simply to the ability to select more diverse questions than a likelihood-based approach. Additioanlly, this reaffirms our observation that VLMs are poorly suited for this task off-the-shelf.
However, we also see that in most cases, Diversity-based ranking substantially degrades relevance, informativeness, number of iterations, and information gain compared to those in Table 2. This suggests that the highly exploratory nature of diversity-based ranking causes rationales to be less coherent, and conclusions are made slower and less confidently with this approach. Meanwhile, coherence-based ranking enables us to find good questions to ask faster, leading to more confident conclusions with more relevant and informative supporting evidence (while also achieving a higher accuracy).
InstructBLIP
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.1 0.2 0.25
10 0.1 0.1 0.23
LLaVA
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.05 0.05 0.58
10 0.05 0.05 0.69
Llama 3
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.05 0.05 0.41
10 0.1 0.025 0.40
Table 15: Selected inference hyperparameters for diversity-based ranking with various VLMs, with optional supplementary candidates generated through in-context learning (ICL).
InstructBLIP
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 61.4 17.8 .244 2.89 .268
Validation 64.0 14.4 .337 3.82 .350
Testing 64.2 17.5 .237 2.88 .268
Testing 64.7 15.1 .337 3.72 .344
LLaVA
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 64.2 37.0 .298 4.12 .444
Validation 65.2 43.2 .385 3.99 .539
Testing 62.9 35.2 .310 4.35 .455
Testing 67.1 39.8 .405 4.32 .513
Llama 3
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 62.6 15.7 .313 6.19 .262
Validation 66.8 24.2 .412 5.14 .358
Testing 60.9 15.1 .322 6.48 .246
Testing 59.6 22.6 .393 5.23 .332
Table 16: Ego4D-PMD validation and testing set results for diversity-based ranking with various VLMs, with optional supplementary candidates generated through in-context learning (ICL).
E.9 DPO with Length Penalty Experiment
In the results discussed in Section 5.2, we observed that while VLMs learned to generate much more relevant questions, the informativeness of answers and thus the PMD accuracy degraded. We hypothesized that this resulted from the generation of highly complex questions, e.g., “Is the soil placed around the seedling with the trowel in the person’s hand?” As such, a potentially fruitful avenue for future research is to explore decoding approaches and learning objectives that prioritize more approachable questions for VLMs.
While the primary purpose of this work was to recast the problem of PMD into a more transparent formulation and lay a foundation for research toward coherent question generation and answering in PMD, we performed an initial experiment to guide future work along this line. Specifically, we imposed an exponential length penalty $l=-1.0$ to text generation during training data generation and inference for evaluation. [^26] Like the results presented in Table 2, candidates generated through in-context learning are also included in training data. During beam search, a length penalty modifies the total log-likelihood $p_{Q}$ of a partial candidate question $Q$ as follows:
$p^{\prime}_{Q}=\frac{p_{Q}}{{|Q|}^{l}}$
Here, $|Q|$ is the length of the question $Q$, i.e., number of generated tokens thus far. For earlier experiments, the value of $l$ could be thought of as 1, which effectively applies no exponential penalty to $|Q|$ when calculating likelihood-based sequence scores (the default behavior).
In Table 17, we list the results of applying this length penalty. As shown, compared to the results in Table 2, similar trends of hold despite applying the length penalty: we observe improved relevance, number of iterations, and information gain, but degraded accuracy and informativeness. An interesting exception is when using likelihood-based ranking and in-context learning during inference, we recover a comparable accuracy and informativeness to those observed in LLaVA before applying DPO, but this comes at a cost of a a lower relevance, higher number of iterations, and lower information gain than other inference configurations. This provides further evidence that there exists a trade-off between generating relevant questions and achieving high informativeness and accuracy, and future work should aim to find a balance between these priorities.
LLaVA + DPO (with length penalty)
Rank ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
L 61.7 93.5 .293 1.93 .741
L 67.0 56.4 .459 4.05 .628
C 63.1 97.5 .313 1.82 .771
C 62.9 97.7 .319 1.83 .769
Table 17: Ego4D-PMD test set results for DPO-trained VLMs with an additional length penalty $l=-1$ applied during training data generation and inference. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.10 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs in Figure 4, in Figure 7, we provide several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.
Correct and coherent points.
Cyan points have low error with high informativeness and relevance, indicating correct decisions with coherent rationales. These are the best case examples from the model. Figure 7, Example A is one such case, where LLaVA correctly determines that the procedure “Pick up a sink brush from the kitchen slab” has been successfully completed, rationalizing it coherently and succinctly with a single question and answer about the location of the sink brush.
Incorrect and incoherent points.
Conversely, red to magenta points have high error, low informativeness, and low relevance, indicating incorrect decisions with incoherent rationales. These are the worst case examples from the model. Figure 7, Example B is one such case, where LLaVA incorrectly decides that the procedure “Tighten the screw” was not successfully completed due to the person in the image not wearing various protective gear, an incoherent rationale for the decision.
Correct but incoherent points.
Indigo to black points have low error, but low relevance and informativeness, indicating correct decisions without sufficient rationale. Figure 7, Example C, is an instance of this, where LLaVA correctly decides that the person in the image has not successfully completed the procedure “Paint the stone” (rather they are painting a wood molding). However, LLaVA’s decision is only supported by a question about whether the person is wearing a shirt, which it does not answer confidently, making for a completely insufficient rationale.
Coherent but incorrect points.
White points have high error, relevance, and informativeness, indicating coherent rationales that do not lead to a correct decision. In other words, the information collected by the VLM should have been sufficient to make a correct decision (according to our automated coherence metrics), but this did not occur. Figure 7, Example D shows one such case, where LLaVA incorrectly decides that the procedure “Drop the bottle of mustard on the countertop” was unsuccessful. While it correctly identified that the bottle is on the countertop, which suggests the success of the procedure, it later mistakenly identified the bottle to be on the floor, creating a contradiction in the rationale and causing it to make the wrong decision. The ability of this analysis to easily identify issues like this may be useful for future work in PMD and task guidance, as it enables the detection and thus the correction of bugs in the system’s reasoning.
Irrelevant but informative points.
Blue points have low relevance but relatively high informativeness, indicating irrelevant questions that still yield informative answers. As shown in Figure 7, Example E, this does not necessarily indicate a failure of LLaVA, rather a terse rationale. In this example, LLaVA correctly determines that the procedure “Fold the cut piece of cloth” has not been completed successfully. It reasonably rationalizes this decision by asking about the presence of a piece of cloth and responding with No. The question of whether the person is working with a piece of cloth is deemed somewhat irrelevant by our metrics because if the answer were instead Yes, this would not provide sufficient information to conclude that the procedure was successful. However, since the answer was No, we do have sufficient information to conclude that the procedure is unsuccessful, despite the question being relatively indirect. Blue points may thus point to sufficient rationales which lack some detail or specificity, which are not necessarily problematic to system performance.
Relevant but uninformative points.
Green and yellow points have high relevance but low informativeness, indicating a failure to extract useful information in VQA. Green points have close to zero informativeness, indicating unsure responses in VQA. In Figure 7, Example F, LLaVA rationalizes its decision about the procedure “Put the bolt remover in the lawn tractor” by asking whether the bolt remover is in the lawn tractor in various ways. However, these objects are not present in the image and thus LLaVA’s answer is not confident, causing it to respond Unsure to most questions, leading to low informativeness. Despite its failure to answer questions, LLaVA still arrives at the correct conclusion that the procedure has not been successfully completed.
Meanwhile, yellow points have highly negative informativeness, indicating counterproductive responses in VQA that oppose the correct decision. As shown in Figure 7, Examples G and H, these cases typically occur when the VLM does not recognize an object in the image, or it recognizes an object that is not in the image. In Example G, LLaVA incorrectly decides that the procedure “Put the trowel in a bin” is unsuccessful because it does not recognize that the trowel is indeed in a bin, perhaps because it is relatively small in the image and does not contrast from the background. In Example H, LLaVA incorrectly decides that the procedure “Put the bottle in the cabinet” is successful because it hallucinates that a bottle is in the cabinet, despite neither object appearing in the image. The ability of this analysis to easily identify failures of visual perception in VLMs again may be useful for future work in this area.
The colors of these points can be used to characterize common behaviors of VLMs. Additional insights toward the fine-grained strengths and weaknesses of various approaches may be gained from analyzing these results by mistake type, or verb and noun categories.
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label: ✓        Predicted: ✓ Rationale: 1. Is the sink brush in the person’s hand? Yes
Example B: Tighten the screw. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the person wearing gloves? No 2. Is the person wearing protective gear? No 3. Is the person wearing a mask? No
Example C: Paint the stone. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person wearing a shirt? Unsure
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the bottle of mustard on the countertop? Yes 2. Is the bottle of mustard on the floor? Yes
Example E: Fold the cut piece of cloth. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person working on a piece of cloth? No
Example F: Put the bolt remover in the lawn tractor. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the bolt remover in the lawn tractor? Unsure 2. Is the bolt remover inside the lawn tractor? Unsure 3. Is the bolt remover attached to the lawn tractor? Unsure
Example G: Put the trowel in a bin. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the trowel in a bin? No
Example H: Put the bottle in the cabinet. Refer to caption Label: ✗        Predicted: ✓ Rationale: 1. Is the bottle in the cabinet? Yes
Figure 7: Sample outputs from LLaVA with coherence-based ranking, representing the range of behaviors visualized in Figure 4. Images cropped for visibility and space.